### v2

# Dirichlet energy and focusing NLS condensates of minimal intensity

Abstract
We consider the family of (poly)continua $\mathcal{K}$ in the upper half-plane $\mathbb{H}$}} that contain a preassigned finite anchor set $E\in{\mathbb{H}}$. For a given harmonic external field we define a Dirichlet energy functional $\mathcal{I}(\mathcal{K})$ and show that within each “connectivity class” of the family, there exists a minimizing compact $\mathcal{K}^{*}$ consisting of critical trajectories of a quadratic differential. In many cases this quadratic differential coincides with the square of the real normalized quasimomentum differential ${\mathrm{d}}\mathbf{p}$ associated with the finite gap solutions of the focusing Nonlinear Schrödinger equation (fNLS) defined by a hyperelliptic Riemann surface $\mathfrak{R}$ branched at the points $E\cup\bar{E}$.
The motivation for this work lies in the problem of soliton condensate of least average intensity such that a given anchor set $E$ belongs to the poly-continuum $\mathcal{K}$. An fNLS soliton condensate is defined by a compact $\mathcal{K}\subset{{\mathbb{H}}}$ (its spectral support) whereas the average intensity of the condensate is proportional to $\mathcal{I}(\mathcal{K})$. We prove that the spectral support $\mathcal{K}^{*}$ provides the fNLS soliton condensate of the least average intensity within a given “connectivity class”.
1 Introduction
A continuum is a compact, connected set with at least two distinct points and a poly-continuum is a finite union thereof. Let $E\subset{\mathbb{C}}$ be a finite set of points, called “anchors”, and $\mathcal{K}\subset{\mathbb{C}}$ be a continuum containing $E$. The well known Chebotarev’s continuum problem [29] is to find such a continuum $\mathcal{K}$ of minimal logarithmic capacity ${\rm cap}(\mathcal{K})$. We recall that ${\rm cap}(\mathcal{K}):={\rm e}^{-\mathcal{E}(\mathcal{K})}$, where
$\mathcal{E}(\mathcal{K}):=\inf\left\{\iint\ln\frac{1}{|z-w|}{\mathrm{d}}\mu(z)d\mu(w)\right\}$ (1.1)
taken among all positive unit Borel measures supported within $\mathcal{K}$. Thus the problem can be stated as that of finding the maximizer of the logarithmic energy $\mathcal{E}(\mathcal{K})$ among all the continua containing $E$. This problem was solved around 1930 in [13], [24, 25], see for example [26], where the minimizing ${\rm cap}(\mathcal{K})$ compact $\mathcal{K}$ was represented as the set of critical trajectories of a certain quadratic differential on the hyperelliptic Riemann surface defined by $E$.
The main problem considered in this paper is of a similar nature but with a different energy functional, a different class of measures involved and a bit different overall setting. Namely, the set $E$ belongs to the upper half plane ${{\mathbb{H}}}=\{z\in{\mathbb{C}}:\ \mathrm{Im}\,(z)>0\}$ and instead of the free logarithmic energy (1.1) we consider the Green energy
$\mathfrak{J}(\mathcal{K}):=$ $\inf_{{\mathrm{d}}\mu}J_{0}[{\mathrm{d}}\mu]$ (1.2)
$J_{0}[{\mathrm{d}}\mu]:=$ $\iint\ln\left|\frac{z-\overline{w}}{z-w}\right|{\mathrm{d}}\mu(z){\mathrm{d}}\mu(w)-2\!\!\int\!\!\mathrm{Im}\,(z){\mathrm{d}}\mu(z)$ (1.3)
with the infimum taken over all nonnegative (but of arbitrary total mass) Borel measures supported on $\mathcal{K}\subset\overline{{{\mathbb{H}}}}$. Here $2\mathrm{Im}\,z$ represents the external field, so (1.2) represents the weighted Green energy $\mathfrak{J}(\mathcal{K})$ of the minimizing measure on $\mathcal{K}$: observe that since $\mathcal{K}\subset\overline{{{\mathbb{H}}}}$, we have $\mathfrak{J}(\mathcal{K})<0$, because the minimum must be smaller than the value for the zero measure. Our extremal problem is thus the one of maximizing the Green’s energy $\mathfrak{J}$ over appropriate classes of (poly-)continua containing the set of anchors $E$. More specifically, we will be looking at the extremal problem of maximizing $\mathfrak{J}(\mathcal{K})$ not only in the class of continua, but in the larger class consisting of poly-continua $\mathcal{K}\supset E$, where each connected component of $\mathcal{K}$ contains at least two points of $E$, or connects a point of $E$ with $\mathbb{R}$}. To the best of our knowledge, this type of extremal problems were not considered in the literature. The most relevant statement we could find is Theorem 6.1 from [31], stated without full proof, where $\mathfrak{J}$ is a Green energy functional on positive Borel measures of total mass one. To give an immediate visual example of some such continua, see Fig. 1, depicting some continua of maximal weighted Green energy (1.2) with the property that $\mathcal{K}\supset E$ and $\mathcal{K}\cap{\mathbb{R}}\neq\emptyset$.
Refer to captionRefer to caption
Figure 1: Various examples of minimal energy sets. These are also examples of solutions of the generalized Chebotarev problem discussed in Problem B.1.
The interest in this extremal (max-min) problem originates from the problem of finding a compact spectral support, $\Gamma^{+}$, of the fNLS soliton condensate of minimal average intensity, given a fixed finite “anchor” set $E\in\Gamma^{+}$. Similar problem can be formulated about the Zakharov–Shabat (continuous) spectrum of finite gap fNLS solutions defined by spectral hyperelliptic Riemann surface branched at $\hat{E}=E\cup\bar{E}$.
In the next Subsection 1.1 we provide a detailed description of the results of the paper, while in the remaining subsections 1.2 of this introduction we discuss the connection and motivation for studying the problem coming from the theory of fNLS soliton condensates.
1.1 Detailed description of the results
In order to formulate the main theorem we need to define the notion of connectivity associated with a poly-continuum. Given a finite anchor set $E\subset{{\mathbb{H}}}$ we define the family ${\mathbb{K}}_{E}$ consisting of all poly-continua $\mathcal{K}$, where each component contains at least two different anchor points or connects an anchor $e\in E$ with a point in $\mathbb{R}$}. We denote the components as $\mathcal{K}=\sqcup_{\ell=0}^{k}\mathcal{K}_{\ell}$, where $\mathcal{K}_{\ell}$, $\ell=1,\dots,k,$ are the connected components of $\mathcal{K}$ not meeting $\mathbb{R}$}, and the notation $\mathcal{K}_{0}$ is reserved for the component of $\mathcal{K}\cup{\mathbb{R}}$ containing $\mathbb{R}$}. This partition of a poly-continuum $\mathcal{K}\in{\mathbb{K}}_{E}$ allow us to define the connectivity of $\mathcal{K}$ as an $(N+1)$ by $(N+1)$ symmetric matrix $M=M(\mathcal{K})$ (connectivity matrix) as follows: the entry $M_{i,j}=1$ means that $e_{i},e_{j}$ belong to the same component $\mathcal{K}_{\ell}$, $0\leq\ell\leq k$,   $1\leq i,j\leq N$, and $M_{i,0}=1$ means that $e_{i}$ belongs to $\mathcal{K}_{0}$; the remaining entries of $M$ are zeros. We say that poly-continua $\mathcal{K}^{(1)},\mathcal{K}^{(2)}\in{\mathbb{K}}_{E}$ have the same connectivity if $M(\mathcal{K}^{(1)})=M(\mathcal{K}^{(2)})$. We say that $\mathcal{K}^{(2)}$ has a connectivity that is (weakly) greater than (or exceeds the connectivity of) $\mathcal{K}^{(1)}$, if $M(\mathcal{K}^{(1)})_{i,j}\leq M(\mathcal{K}^{(2)})_{i,j}$, $i,j=0,\dots,N$ and we write $M(\mathcal{K}^{(1)})\preceq M(\mathcal{K}^{(2)})$; this provides a partial order in the set of allowed connectivities. It is clear that there are finitely many different connectivities within the class ${\mathbb{K}}_{E}$. For a given anchor set $E$ we define subclasses ${\mathbb{K}}_{E,M}\subset{\mathbb{K}}_{E}$, where each subclass consists of poly-continua with larger connectivity than the one defined by the matrix $M$. We can now formulate our main results.
We consider the Dirichlet energy functional, $\mathcal{I}$, defined as follows (see Section 3): let $G(z)$ be the unique solution of the following Dirichlet problem on ${\mathbb{H}}\setminus\mathcal{K}$, namely the function which is harmonic on ${\mathbb{H}}\setminus\mathcal{K}$, continuous and bounded on $\mathbb{H}$$\cup{\mathbb{R}}$, and satisfies $G(z)=\mathrm{Im}\,z$ for $z\in\mathcal{K}\cup{\mathbb{R}}$. The existence is guaranteed under the assumption that $\mathcal{K}$ is a poly-continuum (with finitely many components) by standard results in harmonic analysis, see Section 3. Then the Dirichlet energy of $\mathcal{K}$ is defined by the integral (the factor $\pi$ is here for convenience)
$\mathcal{I}(\mathcal{K}):=\frac{1}{\pi}\iint_{{\mathbb{H}}}\left((\partial_{x}G)^{2}+(\partial_{y}G)^{2}\right){\mathrm{d}}x{\mathrm{d}}y.$ (1.4)
This is a strictly positive quantity. In fact (Section 3) the Dirichlet energy and the Green’s energy $\mathfrak{J}$ are simply related:
$\mathcal{I}(\mathcal{K})=-2\mathfrak{J}(\mathcal{K}).$ (1.5)
so that minimizing one is the same as maximizing the other. See discussion leading to (3.4). We show in Section 5 that the Dirichlet energy is continuous in Hausdorff topology over the class $\mathbb{K}_{E}$ (we actually show a more general result of continuity). See Problem 5.1 and Theorem 5.6.
Existence of a minimizer.
The class $\mathbb{K}_{E}$ and each of the sub-classes $\mathbb{K}_{E,M}$ with preassigned minimal connectivity are all closed in Hausdorff topology. We are interested in the issue of existence (and possibly uniqueness) of the minimizer of $\mathcal{I}$ on these subclasses: if they were compact (alas, they are not), then the existence would be a triviality given the established continuity. Thus we can formulate the first result as follows:
Theorem 1.1
For every connectivity $M$ the Dirichlet energy functional $\mathcal{I}:\mathbb{K}_{E,M}\to{\mathbb{R}}_{+}$ attains a minimum value at a poly-continuum $\mathfrak{F}\in\mathbb{K}_{E,M}$.
Note that no mention is made of the uniqueness, which in general we cannot guarantee. In order to prove Theorem 1.1, we use the fact that each class ${\mathbb{K}}_{E,M}$ is closed in Hausdorff (metric) topology and then we prove that the energy functional $\mathcal{I}(\mathcal{K})$ is continuous in that topology. To prove the existence of a minimizer one needs to show that any minimizing sequence of poly-continua in $\mathbb{K}_{E,M}$ is uniformly bounded. The proof of this theorem consists of Section 5 establishing the continuity of the Dirichlet energy, and Section 6, where the uniform boundedness is shown. This is established by first guaranteeing that the minimizing sequence remains in a horizontal strip (Lemma 6.1) and then that the minimizing sequence is also bounded on the left and right (Lemma 6.2).
A key tool in the proof is provided by the comparison theorem established in Section 4, where we define the Jenkins’ interception property, a generalization of an idea of Jenkins’ formulated in [17]. The proof of this property is based on the “length-area” method [13]. We make some extension in the original statement of this method [16], so that it allows us to compare Dirichlet energies of different $\mathcal{K}\in{\mathbb{K}}_{E}$, provided that certain “interception conditions” are met, see Definition 4.1.
Geometry of the minimizers.
The next question is to characterize the geometrical properties of a minimizer $\mathfrak{F}$ in a given class. We need to introduce the notion of Boutroux quadratic differential of quasi-momentum type (BM for short). This is a quadratic differential $Q(z){\mathrm{d}}z^{2}$ where $Q(z)$ is of the form
$Q(z){\mathrm{d}}z^{2}=\frac{P_{2N}(z)}{\prod_{j=1}^{N}(z-e_{j})(z-\bar{e}_{j})}{\mathrm{d}}z^{2},\quad\text{where}\quad P_{2N}=z^{2N}-2\left(\sum_{j=1}^{N}\mathrm{Re}\,e_{j}\right)z^{2N-1}+\dots\,$, (1.6)
with $P(z)$ a polynomial with real coefficients and such that $Q$ additionally satisfies the Boutroux condition: this means that all contour integrals of $w=\sqrt{Q(z)}$ along closed loops $\gamma$ on the Riemann surface of $w^{2}=Q(z)$ (hyperelliptic) are purely real, i.e.
$\oint_{\gamma}\sqrt{Q(z)}{\mathrm{d}}z\in{\mathbb{R}}.$ (1.7)
The vanishing of the imaginary part of (1.7) gives sufficient real implicit conditions for the coefficients of $P_{2N}$ so that there are finitely many solutions with any given set of anchors $E$. There is an interesting conformal geometry associated to this notion for which we refer to the classical literature [16, 35]. Here suffices to note that the function
$V(z):=\left|\mathrm{Im}\,\int_{0}^{z}\sqrt{Q(\zeta)}{\mathrm{d}}\zeta\right|,$ (1.8)
is well–defined (i.e. independent of the path of integration), harmonic away from its zero level set and behaves near $z=\infty$ like
$V(z)=\mathrm{Im}\,z+I_{Q}\mathrm{Im}\,\left(\frac{1}{z}\right)+\mathcal{O}(|z|^{-2}),$ (1.9)
where the constant $I_{Q}$ is a quantity of interest as we presently see. We denote the zero-level set of $V$ as
$\mathfrak{F}_{Q}:=V^{-1}\big(\{0\}\big).$ (1.10)
We then have
Theorem 1.2
Any minimizer, $\mathfrak{F}$, of the Dirichlet energy within a class $\mathbb{K}_{E,M}$ consists of the zero level curves of the function $V(z)$ (1.8) associated with a suitable Boutroux quadratic differential, $Q$, of quasi-momentum type (1.6). Furthermore the minimum energy $\mathcal{I}(\mathfrak{F}_{Q})$ equals the coefficient $I_{Q}$ in the expansion (1.9).
This type of results is certainly not surprising as this is but a manifestation of the $S$-property [31, 34], which we discuss and recall at the end of Section 2 (see discussion around (2.18)). The Theorem 1.2 is proved by using the Schiffer variation approach, see Section 7.
On the uniqueness of the minimizer.
The final result consists in answering in a partial way the issue of uniqueness. For fixed set of anchors $E$ we choose one of the finitely many Boutroux quadratic differential of quasi-momentum type and $\mathfrak{F}_{Q}=V^{-1}(\{0\})$, see (1.10). This poly-continuum defines a certain connectivity class $M(\mathfrak{F}_{Q})$, where we recall that $M(\mathfrak{F}_{Q})$ denotes the connectivity matrix associated with the poly-continuum $\mathfrak{F}_{Q}$, see previous section. Then
Theorem 1.3
In the class $\mathbb{K}_{E,M({\mathfrak{F}_{Q}})}$ the minimum is unique and is attained at $\mathfrak{F}_{Q}$.
[…]
To belabour the point, we can take Theorem 1.3 as saying that all the local minima of $\mathcal{I}$ over $\mathbb{K}_{E}$ correspond to the set $\mathfrak{F}_{Q}$ associated with a quadratic differential $Q$ as in (1.6).
Electrostatic interpretation.
Let us consider the following two–dimensional electrostatic interpretation of the Dirichlet problem for $V(z)$: we can imagine that a poly-continuum $\mathcal{K}=\cup_{\ell=0}^{k}\mathcal{K}_{\ell}$, see above, is a conductor made of metal with (shielded) connection of each $\mathcal{K}_{l}$, $l\geq 1$, to the ground represented by $\mathbb{R}$}. Namely, we imagine that each $\mathcal{K}_{l}$, $l\geq 1$, is “floating in the sky” and that there is a shielded wire connecting it to the ground, whereas $\mathcal{K}_{0}$, if not empty, consists of grounded conductor(s).
Suppose that we have very high charged clouds (ideally placed at $i\infty$) generating a constant vertical electric field and hence electrostatic potential $\mathrm{Im}\,(z)$; then the conductors $\mathcal{K}$ will distribute charge (of which there is an infinite reservoir via the ground connection) so as to ground themselves at zero potential. The resulting electrostatic potential is precisely $V(z)$ and the Dirichlet energy represents the stored electrostatic energy in this system. Assume further that the conductors $\mathcal{K}_{l}$, $l=0,\dots,k$, can be elastically deformed (with no loss of energy). The restriction is that such deformations should have fixed points from the finite set $E$ assigned to each conductor (according to the connectivity matrix $M$, where $\mathcal{K}\in{\mathbb{K}}_{E,M}$). Then the problem is to find the shape of $\mathcal{K}$ minimizing the electrostatic energy of $\mathcal{K}$.
Remark 1.4
All the pictures in this paper were obtained using a code to explore various configurations of anchors (and also different external fields) produced by one of the authors and available at [2].
1.2 Focusing NLS and extremal problems
With a view to the applications, we explain the proportionality between two quantities:
  • the minimal Dirichlet energy within a class $\mathbb{K}_{E,M}$, established in Theorem 1.1, Theorem 1.3;
  • the average intensity of a finite gap solution of the focusing Nonlinear Schrödinger Equation (fNLS).
The focusing Nonlinear Schrödinger Equation (fNLS) is the PDE
$i\partial_{t}\psi=-\partial_{x}^{2}\psi-2|\psi|^{2}\psi.$ (1.11)
If $\psi=\psi(x,t)\in L^{2}_{loc}$ is a bounded function of $x$, but not necessarily vanishing at $\pm\infty$, one can consider the “average intensity” of $\psi$, namely the limit
$\mathbb{I}(\psi):=\lim_{L\to\infty}\frac{1}{2L}\int_{-L}^{L}|\psi(x,t)|^{2}{\mathrm{d}}x.$ (1.12)
The average intensity $\mathbb{I}(\psi)$ is conserved in time if $\psi(x,t)$ satisfies (1.11) ([1], [9]): this can be easily seen by using the equation (1.11) and integration by parts.
Important classes of solutions to (1.11) are given by the “finite gap solutions” [1], which are quasi-periodic (in $x$ and in $t$) solutions of (1.11) constructed in the seventies [15]. These solutions involve a hyperelliptic Riemann surface $\mathfrak{R}$ of genus $g$ represented as a double cover of the spectral zz–plane (a sort of “Fourier” variable associated to the solution) branched at $2g+2$ points coming in $g+1$ conjugate pairs.
For a finite-gap solution the (continuous) spectrum of the associated linear problem (see Section 2) is the same set $\mathfrak{F}_{Q}$ (1.10) associated with a Boutroux quadratic differential of quasi-momentum type as in Theorem 1.2: this is explained in Section 2. Then the main relationship is that
$\mathbb{I}(\psi)=2\mathcal{I}(\mathfrak{S}),$ (1.13)
where $\mathfrak{S}$ is the continuous spectrum of the associated linear problem, see (2.3). Namely, the average intensity of a finite-gap solution is proportional to the Dirichlet energy of its spectrum. This is shown in Proposition 2.4. In summary, one has the following relation amongst the three quantities, intensity $\mathbb{I}$ (1.12), Dirichlet energy $\mathcal{I}$ (1.4) and Green’s energy $\mathfrak{J}$(1.2):
$\mathbb{I}(\psi)\mathop{=}2\mathcal{I}(\mathfrak{S})\mathop{=}-4\mathfrak{J}(\mathfrak{S}).$ (1.14)
1.3 Motivation: soliton condensates for integrable equations
We now turn our attention to our motivation to study these energetic problems. Recent literature devotes a considerable effort towards the mathematical study of “soliton gases”; the term is used to refer to a couple of approaches that all involve some limiting procedure taken either on special families of $N$-soliton solutions to (1.11) with $N\to\infty$ or on certain meromorphic differentials on $\mathfrak{R}$ of genus $g$, related with the finite gap solutions of (1.11), with $g\to\infty$. In the first case, ideally, one would want to consider some statistical ensembles of infinitely many solitons but in practice, thus far, the state of the art is rather in the direction of choosing a specific $N$-soliton solution, taking the limit $N\to\infty$, and then addressing questions about the behaviour of the limiting solution ([3, 11]). In the second case, while $g\to\infty$, the spectral bands (the branchcuts of $\mathfrak{R}$) are scaled down at a certain rate and in such a way that they fill densely a certain one or two dimensional compact $\Gamma^{+}\subset{{\mathbb{H}}}$, [7], [36]. (In the notations of the present paper, $\Gamma^{+}=\mathcal{K}$.) This limiting procedure is known as thermodynamic limit. The remarkable difference with the previous approach is that here one is primarily interested in some “macroscopic” observable quantities, such as, for example, the effective speed of an element of the soliton gas or the average intensity (of the gas), rather than in reconstruction of particular realizations of the gas, provided that such limiting realization exist. Ideally one would want to calculate large $g$ limits of some statistical characteristics of the finite gap solutions on $\mathfrak{R}$ of genus $g$, such as, for example, the probability distribution of $|\psi|^{2}$, the moments of this distribution, etc. One of the most important macroscopic property of soliton gases with physical relevance is the limiting average intensity $\mathbb{I}(\psi)$, (1.12). Well known results allow to express $\mathbb{I}$ for any finite gap solution [15] and, thus, a suitable description can be obtained in the large genus (thermodynamic) limit, see below. The main thrust for the present work is to find the compact accumulation (spectral support) set $\Gamma^{+}\subset{{\mathbb{H}}}$ of the growing number of small bands (with $E\subset\Gamma^{+}$) that minimizes the average intensity $\mathbb{I}(\psi)$, given by (1.12), for the special type of the fNLS soliton gases, called soliton condensates, see below for more details.
Figure 2: Zero level curves (blue) $\mathfrak{F}_{Q}$ for all possible BMs $Q$ with the set of anchors E=[0.5+2i,0.5+2i,0.96i]E=[-0.5+2i,0.5+2i,0.96i] are shown here. The Dirichlet energies of the left and right cases are the same, approximately $2.7299$, while for the central case the energy is approximately $2.7354$. This is an example of the fact that in $\mathbb{K}_{E}$ there might be not a unique minimizer (in this case there are two). Note that these two sets have “incommensurable” connectivity, namely, there is no connectivity $M$ that precedes both. If the anchor point on the imaginary axis is moved slightly down the connectivity on the left yields the absolute minimizer, while if we move it slightly up the one on the right is the unique minimizer.
Brief introduction to the spectral theory of soliton gases.
The idea of soliton gas for Korteweg de Vries (KdV) equation goes back to 1971 paper [40] of V. Zakharov, where he calculated the effective velocity of a trial soliton propagating on a multi-soliton background. This background modifies the average speed of the (free) trial soliton because of its repeated interactions with the background solitons, each of which can be regarded as an instantaneous shift of the center (aka as phase shift) of the trial soliton. In the modern language, the setting of [40] corresponds to a diluted KdV soliton gas. In order to study a dense KdV soliton gas, a novel approach was suggested by G. El in [5]. This approach is based on styding the finite gap solutions for the KdV, defined by some hyperelliptic Riemann surface $\mathfrak{R}$, where the number $N$ of the bands is growing but the size of the (bounded) bands simultaneously go to zero at a certain exponential (in $N$) rate. Each individual decaying band correspond to a soliton in this limit, but the key thing is the right scaling of the decaying bands, which could be found in an earlier work [38] of S. Venakides. Such a limit is called the thermodynamic limit in [5]. Among the main results of [5] are the so called Nonlinear Dispersion Relations (NDR), which define the continualized limits $u(z):\Gamma^{+}\to\bar{\mathbb{R}}^{+}$ and $v(z):\Gamma^{+}\to\bar{\mathbb{R}}$ of scaled wavenumbers and of scaled frequences respectively of the finite gap KdV solutions in the thermodynamic limit. In spectral theory of soliton gases, $u(z)$ is called the density of states (DOS) and $v(z)$ - the density of fluxes (DOF).
We note that the spectral problem for the KdV is self adjoint so that all the spectral bands of $\mathfrak{R}$ are on $\mathbb{R}$}. Deriving the NDR for a non self adjoint problem, such as, for example, Zakharov-Shabat problem for the fNLS was achieved in [7]. Since the spectral bands of $\mathfrak{R}$ are now in $\mathbb{C}$} (due to Schwarz symmetry, we could restrict our attention to the upper half plane $\mathbb{H}$}} only), the (general) NDR for fNLS soliton gases are complex. For example, the general first NDR for the fNLS soliton gas with a one dimensional accumulation set (an arc) $\Gamma^{+}\subset{{\mathbb{H}}}$ is
$i\int_{\Gamma^{+}}\left[\ln\frac{\bar{w}-z}{w-z}+i\pi\chi_{z}(w)\right]u(w)|dw|+i\sigma(z)u(z)=z+\tilde{u}(z)$ (1.15)
where: $u,\tilde{u}$ are the solitonic and the carrier densities of states (DOS) respectively; $\chi_{z}(w)$ is the indicator function of the (oriented) arc $\Gamma^{+}$ starting at the beginning of $\Gamma^{+}$ and ending at $z\in\Gamma^{+}$, and; $\sigma\in C(\Gamma^{+})$ is nonnegative on $\Gamma^{+}$. The accumulation set $\Gamma^{+}$ is also called a spectral support set for the corresponding soliton gas. Very often ([6], [7]), the imaginary part of (1.15):
$\int_{\Gamma^{+}}\ln\left|\frac{\bar{w}-z}{w-z}\right|u(w)|dw|+\sigma(z)u(z)=\mathrm{Im}\,z,$ (1.16)
defining the solitonic DOS $u(z)$, is called the first NDR. The general second NDR for the fNLS soliton gas with the same $\Gamma^{+}\subset{{\mathbb{H}}}$ as above has the form
$i\int_{\Gamma^{+}}\left[\ln\frac{\bar{w}-z}{w-z}+i\pi\chi_{z}(w)\right]v(w)|dw|+i\sigma(z)v(z)=-2z^{2}+\tilde{v}(z),$ (1.17)
where $v,\tilde{v}$ are the solitonic and the carrier densities of fluxes (DOF) respectively.
Existence and uniqueness of solution $u(z)\geq 0$ to the first NDR (1.16) with a compact $\Gamma^{+}\subset{{\mathbb{H}}}$ was established [20], subject to some mild restrictions on $\Gamma^{+}$ and $\sigma(z)$. The idea of the proof there was to minimize quadratic energy functional $J_{\sigma}[d\mu]:=J_{0}[d\mu]+\int_{\Gamma^{+}}\sigma ud\mu$ among all non negative Borel measures $\mu$, where $d\mu=u(z)|dz|$. In the special case $\sigma\equiv 0$ on $\Gamma^{+}$ the energy $J_{0}$ of the minimizing measure is the Green energy of $\Gamma^{+}\subset{{\mathbb{H}}}$ with the external field $2\mathrm{Im}\,z$, see [32], Chapter 2. This extremal energy is denoted by $\mathfrak{J}(\Gamma^{+})$ in (1.2).
To have another look on the NDR (1.15), (1.17), we remind that the wavenumbers and frequences of finite gap solutions are represented by periods of quasimomentum ${\mathrm{d}}p_{g}$ and quasienergy ${\mathrm{d}}q_{g}$ meromorphic differentials on $\mathfrak{R}$ of genus $g$ that are normalized in such a way that all their periods are real (real normalized or Boutroux differentials). This normalization uniquely defines ${\mathrm{d}}p_{g},{\mathrm{d}}q_{g}$ (their principlal parts at singular points are fixed). If $A$ cycles are properly oriented small loops around the corresponding shrinking bands of $\mathfrak{R}$, then the $A$ and $B$ periods of ${\mathrm{d}}p_{g}$ are called the solitonic and the carrier wavenumbers of the corresponding finite gap solutions, see [7]. Then the general first NDR (1.15) can be viewed as the thermodynamic limit of the Riemann Bilinear relations between ${\mathrm{d}}p_{g}$ and the normalized holomorphic differentials of $\mathfrak{R}$, see [37], [19]. Similarly, one can obtain (1.17) from ${\mathrm{d}}q_{g}$. One can also observe that the kernel of the integral operator in (1.15)-(1.17) was obtained as the thermodynamic limit of the Riemann Period matrix of $\mathfrak{R}$ of the genus $g$ as $g\to\infty$ (subject to some technical restrictions), where the centers of the corresponding bands of $\mathfrak{R}$ approach the values of $w,z\in\Gamma^{+}$ respectively ([36]).
We point out that the fNLS soliton gas with $\sigma\equiv 0$ on $\Gamma^{+}$ defines a special class of soliton gases known as soliton condensates. For soliton condensates, the derivation of the NDR (as the thermodynamic limit of Riemann Bilinear relations) is still in progress (for soliton gases with $\sigma(z)>0$ it was established in [36]). The condensates have a maximizing property, derived in [20]. Namely, it was shown there that if a compact $\Gamma^{+}\subset{{\mathbb{H}}}$ is fixed but $\sigma(z)\geq 0$ is allowed to vary, then
$\mathfrak{J}(\Gamma^{+})=\min J_{\sigma}[d\mu],$ (1.18)
where the minimum is taken among all $\sigma\in C(\Gamma^{+}),\sigma\geq 0$, and all nonnegative Borel measures on $\Gamma^{+}$. Since for the minimizer $\mu_{\sigma}$ of $J_{\sigma}$ we have $J_{\sigma}[d\mu_{\sigma}]=\int_{\Gamma^{+}}\mathrm{Im}\,(z)\,d\mu_{\sigma}(z)$, and the latter expression is propotional to the average intentensity of the condensate ([36]), we conclude that the soliton condensate has the maximal average intensity among all soliton gases with spectral support $\Gamma^{+}$. That observation naturally led to the following question that triggered the work on this paper. Let a compact $\Gamma^{+}\subset{{\mathbb{H}}}$ be a finite collection of piece-wise smooth contours with the fixed endpoints comprising the set $E\subset{{\mathbb{H}}}$. For a given $E$, find $\Gamma^{+}\subset{{\mathbb{H}}}$ that maximizes the Green energy of the soliton condensate defined by $\Gamma^{+}\subset{{\mathbb{H}}}$ with the external field $2\mathrm{Im}\,z$. Equivalently, one can ask of $\Gamma^{+}\subset{{\mathbb{H}}}$ that minimizes the average intensity $\mathbb{I}=\mathbb{I}(\Gamma^{+})$) of the fNLS soliton condensate, defined by $\Gamma^{+}$. As it was discussed above, this problem can be considered as some generalized version of the Chebotarev’s continuum problem.
Acknowledgements.
The authors would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, UK, and Northumbria University, Newcastle, UK, for support and hospitality during the satelite programme “Emergent phenomena in nonlinear dispersive waves”, where work on this paper was undertaken. AT was supported by a Simons fellowship for his visit. The authors also gratefully thank Evguenii Rakhmanov and Arno Kuijlaars for helpful discussions. The work of MB was supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) grant RGPIN-2023-04747. The work of AT was supported in part by the NSF Grants DMS-2009647, DMS-2407080 and by Simons Foundation award. During the revision of the work MB was supported by the Royal Society Wolfson Visiting Fellowship at the School of Mathematics, Bristol University.
2 Zakharov–Shabat equation and spectra
The Lax pair associated with the Nonlinear Schrödinger equation (1.11) consists of the pair of ODEs
$i\partial_{x}\Psi(x,t;z)=U(x,t;z)\Psi(x,t;z),\qquad i\partial_{t}\Psi(x,t;z)=V(x,t;z)\Psi(x,t;z),$ (2.1)
$U(x,t;z):=\begin{bmatrix}z&{\psi(x,t)}\\ \overline{\psi(x,t)}&-z\end{bmatrix},\qquad V(x,t;z):=2zU(x,t;z)+\begin{bmatrix}-|\psi|^{2}&-i\psi_{x}\\ -i\overline{\psi_{x}}&|\psi|^{2}\end{bmatrix}$ (2.2)
for the matrix–valued function $\Psi(x,t;z)$. Here $\psi(x,t)$ is a complex–valued function in a suitable class, depending on the problem considered. The compatibility of these two equations requires that $\psi(x,t)$ satisfies (1.11). The tt–dependence of $\psi$ is not important in our discussion.
We are interested in the (continuous) spectrum of the first ODE in (2.1); by definition it is the closure of the set of $z\in{\mathbb{H}}$ for which all the entries of the solutions $\Psi(x,t;z)$ remain bounded as functions of $x$:
$\mathfrak{S}(\psi):=\overline{\left\{z\in{\mathbb{H}}:\ \sup_{x\in{\mathbb{R}}}\left|\Psi_{ij}(x,t;z)\right|<\infty\right\}}.$ (2.3)
If $\psi(x,0)$ is a finite gap solution [1] then $\Psi(x,0;z)$ can be written explicitly in terms of Riemann Theta functions associated to a hyperelliptic Riemann surface $\mathfrak{R}_{L}$ of finite genus L1L-1 branched at $2L$ points $\{b_{1},\dots b_{L},\overline{b}_{1},\dots,\overline{b}_{L}\},\ b_{j}\in{{\mathbb{H}}}\setminus{\mathbb{R}}$.
We can then describe the continuous spectrum $\mathfrak{S}(\psi)$ rather explicitly in this case: indeed the general structure of the formula (see [15]) is
$\Psi(x,t;z)=W_{L}(z;x,t){\rm e}^{i{(x\mathbf{p}(z)+t\mathbf{q}(z))}\sigma_{3}},\quad$ (2.4)
where $W_{L}$ is a matrix constructed in terms of Riemann theta functions of the Riemann surface $\mathfrak{R}_{L}$, and $\mathbf{p},\mathbf{q}$ are second-kind Abelian integrals normalized along the $A$-cycles. The fundamental but simple observation is that the whole expression is independent of the choice of $A$-cycles of the surface even if each component of the formula (${\mathrm{d}}\mathbf{p},{\mathrm{d}}\mathbf{q}$ and the entries of $W_{L}$) singularly taken does change. If the cycles are chosen to be the loops surrounding the vertical segments $b_{j},\overline{b}_{j}]$, then $W_{L}(z;x,t)$ remains bounded in $x\in{\mathbb{R}}$ for all fixed $z\in{\mathbb{C}}$ (and also for all fixed times $t\in{\mathbb{R}}$). See Remark 2.2 below. With this choice of cycles the function $\mathbf{p}(z)$ from (2.4) is precisely the real–normalized quasimomentum integral, that is, the anti-derivative of the unique differential of the second kind ${\mathrm{d}}\mathbf{p}$ of the form
${\mathrm{d}}\mathbf{p}(z)=\frac{P_{L}(z)}{\sqrt{\prod_{j=1}^{L}(z-b_{j})(z-\overline{b}_{j})}}{\mathrm{d}}z,$ (2.5)
with $P_{L}(z)$ a suitable monic polynomial of degree $L$, uniquely determined by the condition that all the periods of ${\mathrm{d}}\mathbf{p}$ on $\mathfrak{R}_{L}$ are purely real (this is the meaning of real-normalized) and second–kind means that there is no residue at $\infty$:
$\oint_{\gamma}{\mathrm{d}}\mathbf{p}\in{\mathbb{R}},\ \ \ \forall\gamma\in H_{0}(\mathfrak{R}_{L}\setminus\{\infty_{\pm}\}),\ \ \ \mathop{\mathrm{res}}_{\infty_{\pm}}{\mathrm{d}}\mathbf{p}=0.$ (2.6)
Here $H_{0}(\mathfrak{R}_{L}\setminus\{\infty_{\pm}\})$ denotes the homology group of the surface with the two points above infinity, $\infty_{\pm}$, deleted. It also follows from the Schwarz symmetry that all the coefficients of $P_{L}$ are real.
Remark 2.1
The real-normalization of ${\mathrm{d}}\mathbf{p}$ (or ${\mathrm{d}}\mathbf{q}$) is equivalent to requiring that all the AA–periods vanish, with the choice of cycles described above. The Abelian integral of this differential is denoted as $\omega(z)$ at the end of [15]. To see this equivalence we note that since $P_{2N}$ has real coefficients any integral between two complex conjugate branchpoints (vertically) is automatically purely imaginary. So requiring that all periods are real is the same as requiring that the $A$-periods are zero. Since either the $A$-normalization or the “real”-normalization uniquely identify the differential, all else being equal, it follows that the $A$-normalized differential is also real-normalized, if we choose the $A$ cycle as indicated.
Remark 2.2
We want to offer a few more comments about the boundedness of $W_{L}$. The Riemann surface $\mathfrak{R}_{L}$ is a real surface, namely, a Riemann surface with anti-holomorphic involution $\varphi$ sending $z\to\overline{z}$, see [8], Ch. VI. If the $A$ cycles are chosen anti-invariant under the involution $\varphi$ (as we have done) and the $B$-cycles consequently invariant, then one can establish the location of zeros of the Riemann $\Theta$-function in the Jacobian. Furthermore, the vectors of periods of ${\mathrm{d}}\mathbf{p},{\mathrm{d}}\mathbf{q}$ belong to the real-part of the Jacobian (appropriately defined, see loc. cit.). Then, inspection of the entries of $W_{L}$ would reveal that, for all $z\in{\mathbb{H}}\setminus\{b_{1},\dots,b_{L}\}$, all the terms involve evaluations of $\Theta$ along the real part of the Jacobian, where $\Theta$ is periodic, and moreover any denominator involved in the expression contains expressions of $\Theta$ that cannot vanish on the real Jacobian. This guarantees the boundedness with respect to $(x,t)\in{\mathbb{R}}^{2}$ of $W_{L}$.
Remark 2.3
The quasienergy differential ${\mathrm{d}}\mathbf{q}$ appearing in (2.4) is also a second kind real normalized meromorphic differentials with behaviour $\left(-4z+\mathcal{O}(z^{-2})\right){\mathrm{d}}z$ as $z\to\infty$ on the main sheet of $\mathfrak{R}_{L}$, see [9].
The antiderivative of ${\mathrm{d}}\mathbf{p}$ appearing in (2.4) can be computed from any of the end-points, say, $b_{1}$:
$\mathbf{p}(z)=\int_{b_{1}}^{z}{\mathrm{d}}\mathbf{p},$ (2.7)
since the choice only affects an overall constant phase of the solution $\psi$.
From the general structure of the solution (2.4) we observe that the matrix $\Psi$ remains bounded for all $x\in{\mathbb{R}}$ if and only if $\mathrm{Im}\,\mathbf{p}(z)=0$ and thus we have established that the spectrum consists of the zero-level set of $\mathrm{Im}\,\mathbf{p}(z)$:
$\mathfrak{S}(\psi)=\{z\in{\mathbb{H}}:\ \ \mathrm{Im}\,\mathbf{p}(z)=0\}.$ (2.8)
Observe that the equation $\mathrm{Im}\,\mathbf{p}(z)=0$ is meaningful because all the periods of ${\mathrm{d}}\mathbf{p}$ are real and hence the integral in (2.7) is defined only up to overall sign (the choice of determination of the square root) but does not depend on the choice of contour of integration.
It is known [15] that the average intensity (1.12) for such a finite-gap solution $\psi$ is given by [^3] This follows from the use of formula (6.13) and the expression of $\omega(z)$ after (6.14) in loc. cit..
$\mathbb{I}(\psi)=2\mathop{\mathrm{res}}_{z=\infty}z{\mathrm{d}}\mathbf{p}$ (2.9)
The previously stated equality is contained in the next Proposition.
Proposition 2.4
The average intensity $\mathbb{I}(\psi)$ (1.12) for a finite-gap solution coincides with twice the Dirichlet energy (1.4) of its spectrum (2.8):
$\mathbb{I}(\psi)=2\mathcal{I}(\mathfrak{S}(\psi)).$ (2.10)
Proof. Consider the function $V(z)=\left|\mathrm{Im}\,\mathbf{p}(z)\right|$. By the observation after (2.8) this is a nonnegative and continuous function on $\mathbb{H}$$\cup{\mathbb{R}}$; it is also harmonic away from the spectrum, namely away from the zero level set of $\mathrm{Im}\,\mathbf{p}(z)$. Since ${\mathrm{d}}\mathbf{p}$ has no residue at infinity we conclude that $\mathbf{p}(z)$ is single-valued on the domain $|z|>R$ for $R$ sufficiently large, where it behaves as $\mathbf{p}(z)\simeq z+\mathcal{O}(1)$. Thus the spectrum must be bounded.
Let us define $\partial:=\frac{1}{2}(\partial_{x}-i\partial_{y})$ (the Wirtinger operator); we then observe that (a simple consequence of the Cauchy-Riemann equations)
${\mathrm{d}}\mathbf{p}=2i\partial V(z){\mathrm{d}}z,\ \ z\in{{\mathbb{H}}}\setminus\mathfrak{F}.$ (2.11)
This formula is equivalent to requiring that the branch-cut of the square root in ${\mathrm{d}}\mathbf{p}$ be chosen to coincide with $\mathfrak{S}$. The spectrum $\mathfrak{S}$ consists of finitely many analytic arcs that are horizontal trajectories of the quadratic differential $({\mathrm{d}}\mathbf{p})^{2}$ [35], containing the anchors as endpoints. In particular if arcs meet they do so transversally at one of the zeros of the quadratic differential $({\mathrm{d}}\mathbf{p})^{2}$ forming relative angles $\frac{2\pi}{\mu+2}$, with $\mu$ the multiplicity of the zero.
Now define $G(z):=\mathrm{Im}\,z-V$; this is a bounded continuous function on $\mathbb{H}$$\cup{\mathbb{R}}$ and harmonic away from $\mathfrak{S}$. Over $\mathbb{C}$} the function $G$ (as well as $V$) satisfy the Schwarz symmetry
$G(\overline{z})=-G(z).$
Since $V\big|_{\mathfrak{S}}=0$ we have $G=\mathrm{Im}\,z$ on the spectrum $\mathfrak{S}$. Thus, by the minimum principle, $G$ must be positive over all $\mathbb{H}$} since the boundary values are 0 on $\mathbb{R}$} and $\mathrm{Im}\,z>0$ on the boundary of its domain of harmonicity. This shows that $G$ is the solution of the Dirichlet problem used to define the energy (1.4).
We then proceed to evaluate the average intensity using Its-Kotlyarov’s formula (2.9) and (2.11):
$\mathbb{I}(\psi)=\oint_{\gamma}z\frac{{\mathrm{d}}\mathbf{p}}{i\pi}=\frac{2}{\pi}\mathrm{Im}\,\oint_{\gamma^{+}}2iz\partial V(z){\mathrm{d}}z,$ (2.12)
where the integration contour $\gamma$ is a circle in the clockwise direction containing $\mathfrak{S}$ and $\gamma^{+}$ is a (Schwarz symmetrical) deformation of $\gamma$ in $\mathbb{H}$$\cup{\mathbb{R}}$ so that it encircles $\mathfrak{S}$. ($\gamma^{+}$ should contain all points of $\mathfrak{S}\cap{\mathbb{R}}$ and may consist of several loops.) The latter equation follows from the fact that $\frac{{\mathrm{d}}\mathbf{p}}{{\mathrm{d}}z}$ is Schwarz symmetrical. Let us choose an orientation for each smooth arc of $\mathfrak{S}$ (the final formula will not depend on the choice). Let $\theta(z)$ denote the argument of the tangent vector ${\mathrm{d}}z$ along the oriented contour $\mathfrak{S}$. Let $z_{\pm}$ denote the non-tangential boundary values on the left (++) and right (-) of each arc, and let $\partial_{t},\partial_{n_{\pm}}$ denote the tangential directional derivative along $\mathfrak{S}$ and the left/right normal derivatives respectively. Then we have
${\mathrm{d}}z={\rm e}^{i\theta(z)}|{\mathrm{d}}z|,\ \ \ \ 2\partial V(z_{+})={\rm e}^{-i\theta(z)}\left(\partial_{t}-i\partial_{n_{+}}\right)V(z_{+}),\ \ 2\partial V(z_{-})={\rm e}^{-i\theta(z)}\left(\partial_{t}+i\partial_{n_{-}}\right)V(z_{-})\ .$ (2.13)
Thus, using also that $\partial_{n_{+}}=-\partial_{n_{-}}$ on account of the fact that the normals are opposite, we have
$\mathbb{I}(\psi)=$ $\frac{2}{\pi}\mathrm{Im}\,\int_{\mathfrak{S}}iz(\partial_{x}-i\partial_{y})\left(V(z_{+})-V(z_{-})\right){{\mathrm{d}}z}=\frac{2}{\pi}\mathrm{Im}\,\int_{\mathfrak{S}}iz(\partial_{t}-i\partial_{n_{+}})\left(V(z_{+})-V(z_{-})\right){|{\mathrm{d}}z|}=$
$=\frac{2}{\pi}\int_{\mathfrak{S}}\mathrm{Im}\,z\left(\partial_{n_{+}}V(z_{+})+\partial_{n_{-}}V(z_{-})\right){|{\mathrm{d}}z|}=-\frac{2}{\pi}\int_{\mathfrak{S}}G(z)\left(\partial_{n_{+}}G(z_{+})+\partial_{n_{-}}G(z_{-})\right){|{\mathrm{d}}z|},$ (2.14)
In the second to last equation we used the fact that $\mathfrak{S}$ is a zero level set of $V(z)$ so that both boundary values of the tangential derivatives vanish. Thus we have established
$\mathbb{I}(\psi)=\frac{2}{\pi}\int_{\mathbb{H}}\|{\rm grad}\,G(z)\|^{2}{\mathrm{d}}x{\mathrm{d}}y,$ (2.15)
where last equality follows from (2) and the first Green identity. The result is established. $\blacksquare$
2.1 Boutroux Quadratic differentials of quasi-momentum type and finite-gap solutions
To close the circle of identifications we need to tie the discussion of finite-gap solutions with the main problem which is the object of Theorems 1.1, 1.2, 1.3. The square of any quasi-momentum ${\mathrm{d}}\mathbf{p}$ described in the section above, is, by construction, a BM (1.6). Viceversa the square-root of a BM is the quasi-momentum of a finite gap solution. The only subtlety is here in the identification of the bands.
If we consider a BM solution of our minimization problem for given set of anchors, we have it of the form (1.6); the numerator is in general not a perfect square but can be written
$Q(z){\mathrm{d}}z^{2}=\frac{A(z)^{2}\prod_{j=1}^{K}(z-d_{j})(z-\overline{d}_{j}){\mathrm{d}}z^{2}}{\prod_{j=1}^{N}(z-e_{j})(z-\overline{e}_{j})},\ \ \deg A+K=N.$ (2.16)
On the other hand, the quasi-momentum of a finite-gap solution is of the form (2.5): to match the two one has to identify
$Q(z){\mathrm{d}}z^{2}=\frac{A(z)^{2}\prod_{j=1}^{K}(z-d_{j})^{2}(z-\overline{d}_{j})^{2}{\mathrm{d}}z^{2}}{\prod_{j=1}^{K}(z-d_{j})(z-\overline{d}_{j})\prod_{j=1}^{N}(z-e_{j})(z-\overline{e}_{j})}=\frac{P_{L}^{2}(z)}{{\prod_{j=1}^{L}(z-b_{j})(z-\overline{b}_{j})}}{\mathrm{d}}z^{2}=({\mathrm{d}}\mathbf{p}(z))^{2}.$ (2.17)
Namely we are saying that the band endpoints $\{b_{1},\dots,b_{L}\}=E\cup\{d_{1},\dots,d_{K}\}$ with $L=K+N$ and it just so happens that some of the zeros of the polynomial $P_{L}$ in the numerator of (2.5) coincide with some of the denominators. In other words, any BM can be viewed as the quasi-momentum of a finite-gap solution with endpoints of the bands comprising the anchors as well as any zero of odd multiplicity.
Thus we can summarize this discussion with the Proposition
Proposition 2.5
For any connectivity pattern $M$ the minimizer $\mathfrak{F}$ of the Dirichlet energy (1.4) over $\mathbb{K}_{E,M}$ (to be established in Theorems 1.1, 1.2) is the spectrum of a finite–gap solution of the fNLS equation. The average intensity (1.12) of this solution coincides with twice the Dirichlet energy of this minimizer.
The SS–property of the contours.
The arcs of $\mathfrak{F}_{Q}$, see (1.10), have the so–called SS–property, introduced in [31, 34]. At each point $z$ in the (relative) interior of a smooth arc of $\mathfrak{F}_{Q}$ we have two opposite normal directions which we denote by ${\bf n}_{\pm}(z)$. Then the SS–property is the statement that the two normal derivatives of $V=\mathrm{Im}\,\mathbf{p}(z)$ coincide:
$\frac{\partial}{\partial{\bf n}_{+}}V-\frac{\partial}{\partial{\bf n}_{-}}V=0.$ (2.18)
[…]
Figure 3: Examples of four Zakharov–Shabat spectra for the same configuration of anchor set $E$. Indicated also the stagnation points and the trajectories through them. The additional cuts $\Sigma$, used in Prop. 3.6 (not depicted here), would be arcs of orthogonal trajectories extending from the stagnation points upwards toward the blue line representing $\mathcal{K}$, and downwards up to the real axis.
3 Generalized quasimomenta associated to compact sets
Let $\mathcal{K}\subset\overline{{{\mathbb{H}}}}$ be a poly-continuum. We recall that the “outer domain” of $\mathcal{K}$, denoted ${\mathrm{Out}}(\mathcal{K}){:=\Omega}$, is the (unique) unbounded connected component of the complement of $\mathcal{K}$ in $\mathbb{H}$}} and its complement is called the polynomial convex hull [32] of $\mathcal{K}$, $\mathcal{K}\subset{\mathrm{Hull}}(\mathcal{K})$ in general and ${\mathrm{Hull}}(\mathcal{K})$ is also compact. The “outer boundary” of $\mathcal{K}$ is then the boundary of ${\mathrm{Out}}(\mathcal{K})$ (and of ${\mathrm{Hull}}(\mathcal{K})$), which is a subset, in general, of the boundary of $\mathcal{K}$. Now consider the Dirichlet problem
Problem 3.1
Let $G(z)=G(z;\mathcal{K}):{{\mathbb{H}}}\to{\mathbb{R}}_{+}$ be the unique function satisfying the conditions
  1. 1.
    $G(z)$ is continuous and bounded in $\mathbb{H}$$\cup{\mathbb{R}}$;
  2. 2.
    $G(z)$ is harmonic in ${{\mathbb{H}}}\setminus\mathcal{K}$ ;
  3. 3.
    $G\big|_{\mathcal{K}\cup{\mathbb{R}}}\equiv\mathrm{Im}\,z$.
We observe that any continuum (and so poly-continuum) is regular for the Dirichlet problem, meaning that the Green function of the complement is continuous up to the boundary and is zero therein. This follows from the characterization in [32], Appendix A.2, Theorem 2.1, of regular points in terms of the Wiener condition, which is easily shown to hold at all points of the continuum.
Since $\mathrm{Im}\,(z)$ is harmonic in $\mathbb{H}$}} we have $G\equiv\mathrm{Im}\,(z)$ on ${\mathrm{Hull}}(\mathcal{K})$ (which is larger than $\mathcal{K}$ in general), so that the only important information is contained in the shape of ${\mathrm{Hull}}(\mathcal{K})$. Therefore, without loss of generality, we assume henceforth that $\mathcal{K}={\mathrm{Hull}}(\mathcal{K})$. Let $H$ be a (multi-valued) harmonic conjugate function to $G$ in ${\mathrm{Out}}(\mathcal{K})$. The function $H$ has additive multi-valuedness under the harmonic continuation in ${\mathrm{Out}}(\mathcal{K})$ unless the latter is simply connected, which is in general not the case. By performing some additional branchcuts, $H$ becomes single valued in the complement, see Section 3.1.1.
We are now going to show that $G(z;\mathcal{K})$ can be written as
$G(z;\mathcal{K})=\int_{\partial\mathcal{K}}\ln\left|\frac{z-\overline{w}}{z-w}\right|{\mathrm{d}}\rho_{\mathcal{K}}(w),$ (3.1)
where ${\mathrm{d}}\rho_{\mathcal{K}}$ is the minimizing measure of $J_{0}[{\mathrm{d}}\mu]$ in (1.3) amongst all positive measures supported on $\mathcal{K}$. It is well known ([32]) that the measure is supported on the boundary only (because the external field is harmonic) and the variational equation
$\int_{\partial\mathcal{K}}\ln\left|\frac{z-\overline{w}}{z-w}\right|{\mathrm{d}}\rho_{\mathcal{K}}(w)=\mathrm{Im}\,z$ (3.2)
is, in general, valid quasi-everywhere on the support of ${\mathrm{d}}\rho_{K}$.
But since $\mathrm{Im}\,z>0$ in $\mathbb{H}$} and, moreover, $\mathcal{K}$ is a poly-continuum (therefore regular for the Dirichlet Problem 3.1), the equation (3.2) is in fact valid everywhere on $\partial\mathcal{K}$ (see, for example, [20], Section 1.3). The left hand side of (3.2) is known as the Green potential of ${\mathrm{d}}\rho_{\mathcal{K}}$. This function is harmonic in the complement of $\mathcal{K}$ and satisfies the remaining two conditions in Problem 3.1. Thus, the Green potential of the minimizer ${\mathrm{d}}\rho_{\mathcal{K}}$ solves the Dirichlet Problem 3.1. Moreover the support of ${\mathrm{d}}\rho_{\mathcal{K}}$ coincides with $\partial\mathcal{K}$ ([20], Proposition 3.5) and, according to (3.2), (1.3),
$\mathfrak{J}(\mathcal{K})=\int(G(z;\mathcal{K})-2\mathrm{Im}\,z){\mathrm{d}}\rho_{\mathcal{K}}(z)=-\int\mathrm{Im}\,(z){\mathrm{d}}\rho_{\mathcal{K}}(z).$ (3.3)
We can now use the first Green identity and the well known restoration formula for a measure from its potential [32] to derive [^4] For simplicity we assume sufficient smmothness (piece-wise smoothness) of $\partial\mathcal{K}$.
$-2\mathfrak{J}(\mathcal{K})=2\int_{\partial\mathcal{K}}\mathrm{Im}\,z{\mathrm{d}}\rho_{\mathcal{K}}=2\int_{\partial\mathcal{K}}G(z){\mathrm{d}}\rho_{\mathcal{K}}=-\frac{1}{\pi}\int_{{\mathbb{R}}\cup\partial\mathcal{K}}G(z)\left[\frac{\partial G_{+}(z)}{\partial n_{+}}+\frac{\partial G_{-}(z)}{\partial n_{-}}\right]|{\mathrm{d}}z|=$ (3.4)
$\frac{1}{\pi}\int_{{\mathbb{H}}\setminus\partial\mathcal{K}}\left[G(z)\Delta G(z)+|\nabla G(z)|^{2}\right]dA=\frac{1}{\pi}\int_{\mathbb{H}}|\nabla G(z)|^{2}dA=\mathcal{I}(\mathcal{K}),$~~~~~~~~~~~~~~~~~~~~~~~~~~ (3.5)
where $dA$ is the standard Lebesque measure in ${\mathbb{R}}^{2}$.
Remark 3.2
In [20] it is proved that the fact the support of $\rho_{\mathcal{K}}$ coincides with $\partial\mathcal{K}$ requires that the external potential $\phi$ is positive (as is, in our case). If $\phi$ is not everywhere positive and we minimize (1.2) over positive measures only, the equilibrium measure may have a smaller support.
The case of finite-gap spectra.
In the case when $\mathcal{K}$ is the spectrum $\mathfrak{S}$ of a finite-gap solution as explained in Sec. 2 we have the relation of three quantities: the average intensity $\mathbb{I}(\psi)$, the Dirichlet energy $\mathcal{I}(\mathfrak{S})$ and the Green’s energy:
[…]
In particular we have the equivalence
${\mathcal{I}}(\mathcal{K})=\frac{1}{\pi}\iint_{{\mathbb{H}}}\big|{\rm grad}\,G\big|^{2}{\mathrm{d}}x{\mathrm{d}}y=2\int_{\partial\mathcal{K}}\!\!\!\mathrm{Im}\,(z){\mathrm{d}}\rho_{\mathcal{K}}(z).$ (3.7)
These identities fully close the circle of ideas that motivate our interest in the Dirichlet energy.
3.1 The generalized quasimomentum and the uniformization theorem
If we define the generalized quasimomentum by
$\mathcal{P}(z;\mathcal{K}):=z-g(z)=z-iG(z;\mathcal{K})+H(z;\mathcal{K})=U(z)+iV(z),$ (3.8)
then it has the property that:
  1. 1.
    $\mathcal{P}$ is analytic (multi-valued) in ${\mathrm{Out}}(\mathcal{K})$;
  2. 2.
    $V=\mathrm{Im}\,(\mathcal{P})\equiv 0$ on $\mathcal{K}\cup{\mathbb{R}}$;
  3. 3.
    as $|z|\to\infty$ we have (as a convergent series)
    $\mathcal{P}(z;\mathcal{K})=z+\frac{2\int_{\partial\mathcal{K}}\mathrm{Im}\,(w){\mathrm{d}}\rho_{\mathcal{K}}(w)}{z}+\sum_{j=2}^{\infty}\frac{2}{z^{j}}\int_{\partial\mathcal{K}}\mathrm{Im}\,(w^{j}){\mathrm{d}}\rho_{\mathcal{K}}(w).$ (3.9)
The relationship between the generalized quasimomentum $\mathcal{P}$ and the uniformizing map is elucidated by the following proposition.
Proposition 3.3
Suppose that $\Omega={\mathrm{Out}}(\mathcal{K})$ is simply connected in $\mathbb{H}$}}; then $\mathcal{P}(z;\mathcal{K})$ is the uniformizing map of $\Omega$ to $\mathbb{H}$}} with the normalization condition that $\mathcal{P}(z;\mathcal{K})=z+\mathcal{O}(z^{-1})$ as $z\to\infty$. In particular, there are no points $z\in\Omega$ with $\mathcal{P}^{\prime}(z;\mathcal{K})=0$, i.e., the points with ${\rm grad}\,G(z;\mathcal{K})=0$.
Proof. Let us temporarily denote by $\varphi(z)$ an uniformizing map of the assumed simply-connected domain ${\mathrm{Out}}(\mathcal{K})$ to the upper half-plane (the existence of which is guaranteed by the Riemann uniformization theorem). We can fix it uniquely if we impose $\varphi(\infty)=\infty$ and then normalize (by real–multiplication and addition of a constant) so that $\varphi(z)=z+\mathcal{O}(z^{-1})$. Since $\varphi$ is a uniformizing map, we have $\mathrm{Im}\,(\varphi)\bigg|_{{\mathbb{R}}\cup\partial{\mathrm{Out}}(\mathcal{K})}\!\!\!\!\!\!\!\!\!\equiv 0$. Thus the function $G(z):=\mathrm{Im}\,(z-\varphi(z))$ solves Problem 3.1 in ${\mathrm{Out}}(\mathcal{K})$. We can extend it to be identically equal to $G\equiv\mathrm{Im}\,(z)$ in the ${\mathrm{Hull}}(K)$ and then it must coincide with the solution of the same Problem 3.1. Thus $\mathrm{Im}\,\varphi(z)=\mathrm{Im}\,(z)-G$. The harmonic conjugate function of $G$, denoted by $H$, is uniquely defined (up to additive constant) in ${\mathrm{Out}}(\mathcal{K})$ because of the assumption that ${\mathrm{Out}}(\mathcal{K})$ is simply connected and hence $\varphi(z)=z-iG+H$. Since $G$ is precisely the solution of Problem 3.1, then the equality $\varphi=\mathcal{P}$ follows from (3.8). $\blacksquare$
3.1.1 The non simply connected case
Even if ${{\mathbb{H}}}\setminus\mathcal{K}$ is not simply connected, the map $\xi=\mathcal{P}(z;\mathcal{K})$ can be used to describe a useful univalent function, provided we perform some additional slits.
Proposition 3.4
Let us denote by $\mathcal{K}_{\ell}$, $\ell=1,\dots,k,$ the connected components of $\mathcal{K}$ not meeting $\mathbb{R}$} and by $\mathcal{K}_{0}$ the remaining connected component of $\mathcal{K}\cup{\mathbb{R}}$. Then $\mathcal{P}^{\prime}(z;\mathcal{K})=\frac{{\mathrm{d}}\mathcal{P}(z;\mathcal{K})}{{\mathrm{d}}z}$ has $k$ zeros (“stagnation points”), counted with multiplicity, in ${{\mathbb{H}}}\setminus\mathcal{K}$.
We are only sketching the elementary proof. It can be established using the Argument Principle, i.e., calculating the increment of $\arg\mathcal{P}^{\prime}(z)$ along a contour $\gamma$ that closely follows the boundary of $\Omega={\mathrm{Out}}(\mathcal{K})$ (but also using elementary Morse theory applied to $\mathrm{Im}\,\mathcal{P}$). For example, let $\gamma$ consist of the union of level curves $\mathrm{Im}\,\mathcal{P}(z)=M$ and $\mathrm{Im}\,\mathcal{P}(z)=m$, where $M>0$ is sufficiently large and $m>0$ is sufficiently small. Due to (3.9), the increment of $\arg\mathcal{P}^{\prime}(z)$ along the “semicircle” $\mathrm{Im}\,\mathcal{P}(z)=M$ is very small. Along the level curves of $\mathrm{Im}\,\mathcal{P}(z)$, we have $\arg\mathcal{P}^{\prime}(z)=-\arg{\mathrm{d}}z$. Thus, we have a small increment of $\arg\mathcal{P}^{\prime}(z)$ along the level curve $\mathrm{Im}\,\mathcal{P}(z)=m$ near the remaining part of the outer boundary of $\Omega$. However, the increment of $\arg\mathcal{P}^{\prime}(z)$ along $\mathrm{Im}\,\mathcal{P}(z)=m$ around each of the remaining components of $\mathcal{K}$ is $2\pi$, which yields the desired statement. Similar result was proven in [28], Section 26, for critical points of a Green function of a multiply connected region.
Remark 3.5
The terminology of ”stagnation points” comes from the interpretation of the integral lines of the gradient of $G(z;\mathcal{K})$ as flow-lines of a two–dimensional fluid and the fact that they are points where the gradient flow has a fixed point.
Denote the stagnation points of Proposition 3.4 by $z_{1},\dots,z_{s}$ with multiplicies $m_{1},\dots,m_{s}$, $\sum_{j=1}^{s}m_{j}=k$.
From each $z_{j}$ we take all the arcs of steepest descent trajectories of $\Phi=\mathrm{Im}\,\mathcal{P}$ up to either another stagnation point or $\mathcal{K}\cup{\mathbb{R}}$. Denote by $\Sigma:=\bigcup_{d}\sigma_{d}$ the union of these arcs.
We claim that there are $e:=\sum_{j=1}^{s}(m_{j}+1)$ such arcs in $\Sigma$. Indeed we have taken all the steepest descent trajectories from the points $z_{j}$ (of which there are $(m_{j}+1)$). In the generic (Morse) case there are precisely $2k$ arcs and all $m_{j}=1$. We want to show
Proposition 3.6
The domain $\mathcal{D}:={{\mathbb{H}}}\setminus(\mathcal{K}\cup\Sigma)$ is connected and simply connected and the map $\xi=\mathcal{P}(z;\mathcal{K})$ is univalent on $\mathcal{D}$ and maps it to the upper half $\xi$–plane minus finitely many vertical segments with one endpoint on the real axis.
For example, if all stagnation points are simple, then the system of cuts $\Sigma$ consists of the union of the two arcs of steepest descent from each such point. See caption of Fig. 3. Then, each such pair of arcs is represented in the $\xi$–plane by a pair of vertical segments of the same length. We first prove the claims of connectedness.
Lemma 3.7
The set ${{\mathbb{H}}}\setminus(\mathcal{K}\cup\Sigma)$ is connected and simply connected.
Proof. This is the same as showing that $\mathcal{K}\cup\Sigma$ is a tree (in the sense of graph theory), with set $\mathcal{V}$ of $k+s+1$ vertices consisting of one node associated to each of the connected components $\mathcal{K}_{0},\dots\mathcal{K}_{k}$, and one for each stagnation point $z_{1},\dots,z_{s}$. The set of edges $\mathcal{E}$ consists of the $\sum_{j=1}^{s}(m_{j}+1)=s+k$ arcs in $\Sigma$.
First we show that ${{\mathbb{H}}}\setminus(\mathcal{K}\cup\Sigma)$ is connected, i.e., it has only one connected component; if not, then at least one component, say $\mathscr{D}$, is bounded (since $\Sigma$ is compact and the only non-compact component is $\mathcal{K}_{0}\supset{\mathbb{R}}$). The boundary of such $\mathscr{D}$ consists of pieces of boundaries of various $\mathcal{K}_{\ell}$’s, and edges of $\Sigma$ (and corresponds to a loop in the graph).
The function $\Phi=\mathrm{Im}\,\mathcal{P}$ is harmonic in $\mathscr{D}$ and thus should take its maximum on the boundary of $\mathscr{D}$; this maximum must precisely occur at one of the stagnation points on $\partial\mathscr{D}$ (because $\Phi=0$ on $\mathcal{K}$). Suppose $z_{0}$ is the maximum on the boundary. However $z_{0}$ is a saddle point and the boundary of $\mathscr{D}$ near $z_{0}$ must consist of two branches of steepest descent trajectories [^5] The boundary of $\mathscr{D}$ at a critical point $z_{j}\in\partial\mathscr{D}$ in general could consist also of either two ascending or one ascending and one descending trajectory, but in this case clearly $z_{j}$ would not a local maximum of $\Phi$ along the boundary., which leaves at least one steepest ascent trajectory from $z_{0}$ going into $\mathscr{D}$. So $z_{0}$ could not be a maximum and we have reached the desired contradiction.
Note that since the boundary of $\mathscr{D}$ would constitute a loop in the graph $\mathcal{K}\cup\Sigma$, the above argument shows that $\mathcal{K}\cup\Sigma$ has no loops and hence it is either a tree or a forest (union of disjoint trees). We want to exclude the latter case. This is a simple counting argument with the Euler characteristic: for each tree we must have one less edge than vertices. Since we have a total of $k+s+1$ vertices already and $k+s$ edges, the graph $\mathcal{K}\cup\Sigma$ must be connected. Thus, ${{\mathbb{H}}}\setminus(\mathcal{K}\cup\Sigma)$ must be simply connected. $\blacksquare$
Proof of Proposition 3.6.
The domain $\mathcal{D}:={{\mathbb{H}}}\setminus(\mathcal{K}\cup\Sigma)$ is connected and simply connected and also, by Proposition 3.3, does not contain any zero of $\mathcal{P}^{\prime}$; thus $\mathcal{P}$ can be defined as a univalent and single–valued analytic function. All boundaries of $\mathcal{K}_{0}\supset{\mathbb{R}},\mathcal{K}_{1},\dots,\mathcal{K}_{k}$ are mapped to segments of the real $\xi$–axis, while all the edges of $\Sigma$, by construction being unions of gradient lines of $\mathrm{Im}\,\xi=\mathrm{Im}\,\mathcal{P}$ must have constant $\mathrm{Re}\,\xi=\mathrm{Re}\,\mathcal{P}$ and hence are mapped to vertical slits. Along these slits are the images of stagnation points. In particular, the tip of each slit correspond to a stagnation point. Each arc $\sigma_{h}$ of $\Sigma$ is mapped to a sub-segment of one of the sides of a slit. $\blacksquare$
In the generic case where $\Phi$ is a (nondegenerate) Morse function [27], all the stagnation points are simple, and each of the $2k$ arcs $\sigma_{j}$ connects a stagnation point to one of the components of $\mathcal{K}\cup{\mathbb{R}}$, the image $\mathcal{P}(\mathcal{D})$ is simplest. It consists of $\mathbb{H}$}} with $2k$ pairs of vertical slits with the sides of the slits identified in pairs, and the apex of each pair representing the same stagnation point.
Of course the statements of this section apply equally well to the Zakharov–Shabat spectra; we will however keep the notation $\zeta=\mathbf{p}(z)$ for the uniformizing map in that case. The Zakharov–Shabat case has the additional property of the existence of a measure preserving involution, which is a property ultimately equivalent to Stahl SS–property.
Figure 4: Example of orthogonal flow-lines (level curves of $\mathrm{Re}\,\mathbf{p}(z)$) when $\Omega={\mathrm{Out}}(\mathfrak{F})$ is simply connected. The Zakharov-Shabat spectrum $\mathfrak{F}=\mathfrak{F}_{{}_{ZS}}\cap{{\mathbb{H}}}$ is shown by blue lines, which are zero level curves of $\mathrm{Im}\,\mathbf{p}(z)$. The red points at the end of blue lines form the set $E$. The level curves of $\mathrm{Re}\,\mathbf{p}(z)$ emanated from $\mathfrak{F}_{{}_{ZS}}$ and from $\mathbb{R}$} are shown in light blue and in black respectively. Note the absence of stagnation points in $\mathbb{H}$}}, so that the orthogonal flow from $\mathbb{H}$}} to $\mathfrak{F}\cup{\mathbb{R}}$ is everywhere continuous.
Figure 5: Example of orthogonal flow-lines (level curves of $\mathrm{Re}\,\mathbf{p}(z)$) when $\Omega={\mathrm{Out}}(\mathfrak{F})$ is not simply connected. The Zakharov-Shabat spectrum $\mathfrak{F}=\mathfrak{F}_{{}_{ZS}}\cap{{\mathbb{H}}}$ is shown by blue lines, which are zero level curves of $\mathrm{Im}\,\mathbf{p}(z)$ . The gradient lines of $\mathrm{Im}\,\mathbf{p}(z)$ emanating from $\mathfrak{F}$ and from $\mathbb{R}$} are shown in light blue and in black respectively. Note the stagnation point $z_{0}\in{{\mathbb{H}}}$ (the intersection of green curves) so that the orthogonal flow from $\mathbb{H}$}} to $\mathfrak{F}\cup{\mathbb{R}}$ is discontinuous across the “caustic” (not depicted). The level curve $\mathrm{Im}\,\mathbf{p}(z)=\mathrm{Im}\,\mathbf{p}(z_{0})$ is shown in green.
4 Jenkins’ interception property and the Dirichlet energy
In this section we prove a comparison theorem between the Dirichlet energies of two sets. We will use as reference a set $\mathfrak{F}$ such that ${{\mathbb{H}}}\setminus\mathfrak{F}$ is connected and $\mathfrak{F}$ consists of a finite union of smooth arcs, so that at each relative interior point of each arc we can define the tangent and the two normal directions. We will speak of orthogonal trajectories to mean the integral lines of the gradient of $V$ in (3.8); these are orthogonal to the level-sets of $V$ and they are equivalently described as the trajectories with tangent satisfying $\mathrm{Re}\,{\mathrm{d}}\mathcal{P}=0$. With a slight abuse of notation, we will denote by $\mathbf{p}(z):=\mathcal{P}(z;\mathfrak{F})$ the generalized quasimomentum of $\mathfrak{F}$. In general this set $\mathfrak{F}$ does not have the S-property (2.18), that is, the two normal derivatives of $v=\mathrm{Im}\,\mathbf{p}$ from the opposite sides are not necessarily equal to each other. For each point $z$ in the relative interior of one of the arcs there are exactly two orthogonal flow-lines of the gradient of $v$, and the one in the direction of the larger $\frac{\partial\mathrm{Im}\,\mathbf{p}(z)}{\partial{\bf n}_{\pm}}$, will be called dominant orthogonal trajectory and denoted $\mathcal{L}^{\perp}_{d}(z)$. The remaining orthogonal trajectory will be called recessive and denoted $\mathcal{L}^{\perp}_{r}(z)$. In the case where the normal derivatives are equal, any of the two orthogonal trajectories can be considered as dominant and in this case $\mathcal{L}^{\perp}_{d}(z)$ is the union of both.
Definition 4.1
We say that a poly-continuum $\mathcal{K}$ has Jenkins’ interception property with respect to $\mathfrak{F}$ if $\mathcal{L}^{\perp}_{d}(z)\cap\mathcal{K}\neq\emptyset$ for all $z$ in the relative interior of every arc of $\mathfrak{F}$.
In the case when $\mathfrak{F}$ represents a Zakharov–Shabat spectrum (i.e. it has the $S$-property), the Jenkins’ interception property means that for every $z$ in the interior of $\mathfrak{F}$ at least one of the orthogonal trajectories from $z$ (from one side or the other or both) intersects $\mathcal{K}$. The main results of this paper are based on the following theorem.
Theorem 4.2
If a poly-continuum $\mathcal{K}$ has Jenkins’ interception property with respect to $\mathfrak{F}$ (Def. 4.1) then
${\mathcal{I}}(\mathfrak{F})\leq{\mathcal{I}}(\mathcal{K}).$ (4.1)
Moreover, the equality in (4.1) implies $\mathcal{K}=\mathfrak{F}$.
Theorem 4.2 implies that a small deformation of $\mathfrak{F}$ in the dominant direction increases ${\mathcal{I}}(\mathfrak{F})$. Here “small” means a small and smooth displacement of a point $z$ in the interior of $\mathfrak{F}$ as well as a small variation in the normal direction ${\bf n}(z)$ to $\mathfrak{F}$. In fact, these deformations do not have to be small as long as the connectivity (the topology) of $\mathfrak{F}$ is preserved. The above arguments show that any $\mathfrak{F}$ with S-property is a local minimum of the Dirichlet energy functional ${\mathcal{I}}(\mathfrak{F})$. Moreover, it must be the global minimum of ${\mathcal{I}}(\mathfrak{F})$ in the subclass of compacts in ${\mathbb{K}}_{E}$ with the same connectivity as $\mathfrak{F}$, i.e., in the class of connectivity perserving deformations.
The proof of Theorem 4.2 is based on the “length–area method” [^6] The terminology is used in the literature on Teichmüller theory and appears to originate in the ideas of Grötzsch [13], see historical summary in [39]., and generalizes the result of [17]. For a given $\mathcal{K}$ we have the function $\mathcal{P}(z)=\mathcal{P}(z;\mathcal{K})$ defined as a multi-valued analytic function on ${{\mathbb{H}}}\setminus\mathcal{K}$ with purely real additive multivaluedness, i.e. the analytic continuation along any closed loop in ${{\mathbb{H}}}\setminus\mathcal{K}$ yields the same germ of analytic function plus a real constant (recall our running assumption that $\mathcal{K}={\mathrm{Hull}}(\mathcal{K})$).
From the asymptotics of the quasimomenta (3.9) it follows that both $\mathbf{p}(z),\mathcal{P}(z)$ are invertible for sufficiently large $|z|$.
The proof of Theorem 4.2 is preceeded by several lemmas. We start by denoting by $R_{L}$ the rectangle in the $\zeta=u+iv$–plane of area $2L^{2}$:
$R_{L}:=\Big\{u\in[-L,L],\ v\in[0,L]\Big\}.$ (4.2)
Let us define a deformed rectangle ${\mathcal{R}}_{L}$ in the $z$-plane as the region bounded by the pre-images of the left/right/top sides of $R_{L}$ and the real axis of $\zeta=\mathbf{p}(z)$. Similarly we define the deformed rectangle $\mathscr{R}_{L}$ in the $\xi$–plane, where $\xi=\mathcal{P}(z;\mathcal{K})$. Note that ${\mathcal{R}}_{L}$ in the zz–plane contains all of $\mathfrak{F}$ for $L$ sufficiently large, see Fig. 6.
Figure 6: The three rectangles.
The vertical lines in the $\zeta$–plane correspond to the foliation by the orthogonal flow of $v=\mathrm{Im}\,\mathbf{p}$ in the zz–plane.
We are interested in the single–valued analytic differential $\frac{{\mathrm{d}}\mathcal{P}}{{\mathrm{d}}z}$ on ${{\mathbb{H}}}\setminus\mathcal{K}$. Note that if $z$ is any interior point to $\mathcal{K}$ (in the topological sense), then ${\mathrm{d}}\mathcal{P}/{\mathrm{d}}z=0$ in any small disk centered at $z$ and contained in $\mathcal{K}$.
The conformal area form and metric induced by $\mathcal{P}$ (see [16] for more context) are given by
${\mathrm{d}}A:=\hat{\rho}^{2}{\mathrm{d}}^{2}z,\ \ \ \ \hat{\rho}:=\left|\frac{{\mathrm{d}}\mathcal{P}}{{\mathrm{d}}z}\right|,\ \ \ {\mathrm{d}}^{2}z:={\mathrm{d}}x{\mathrm{d}}y,\ \ {\mathrm{d}}s=\hat{\rho}|{\mathrm{d}}z|.$ (4.3)
Note that in this metric any point in the interior of the same connected component of $\mathcal{K}$ is at zero distance from any other point in the same class.
By the change of variable formula the area form ${\mathrm{d}}A$ in the $z$-plane is transformed to the $\zeta$-plane as follows:
$\mathbf{p}^{\star}({\mathrm{d}}A)=\rho^{2}(u,v){\mathrm{d}}^{2}\zeta,\ \ \ \rho(u,v):=\hat{\rho}\left|\frac{{\mathrm{d}}z}{{\mathrm{d}}\mathbf{p}}\right|=\left|\frac{{\mathrm{d}}\mathcal{P}}{{\mathrm{d}}\mathbf{p}}\right|,\quad\zeta=u+iv,\ \ {\mathrm{d}}^{2}\zeta:={\mathrm{d}}u{\mathrm{d}}v.$ (4.4)
The main strategy is to compute an upper and lower bound for the integral
$\mathscr{A}_{L}:=\iint_{R_{L}}{\rho^{2}(u,v)}{\mathrm{d}}^{2}\zeta,$ (4.5)
which represents area of the image $\mathscr{R}_{L}$ of $R_{L}$ under the map $\mathcal{P}\circ\mathbf{p}^{-1}(R_{L})$.
Lower bound for (4.5).
Let us partition $L,L]=U^{+}\sqcup U^{-}\sqcup U^{0}$ into disjoint sets. To describe them let $\mathfrak{F}^{\circ}$ denote the union of the relative interiors of all arcs of $\mathfrak{F}$ (i.e. all except the meeting points of two or more smooth arcs) and observe that the image $\zeta=\mathbf{p}(z)$ of every point $z\in\mathfrak{F}^{\circ}$ appears exactly twice on [L,L][-L,L], say at $u$ and $u^{\star}$. Our convention is such that $u$ corresponds to the side of the dominant trajectory $\mathcal{L}^{\perp}_{d}(z)$ and we call such $u$’s “dominant”. Viceversa, points in the relative interior of ${\mathbb{R}}\setminus\mathfrak{F}$ appear only once in [L,L][-L,L]. Then we define the dominant subset by
$U^{+}=\left\{u\in[-L,L]:\exists w\in\mathfrak{F}^{\circ}:\ \ \mathbf{p}(w)=u,\text{ $u$ dominant}\right\}.$ (4.6)
The recessive subset is similarly defined
[…]
Finally we denote $U^{0}=[-L,L]\setminus\overline{U^{+}\cup U^{-}}$; this includes the points on $\mathbb{R}$}.
Lemma 4.3
Let $L$ be sufficiently large. For any $u\in U^{+}\cup U^{-}$ we denote by $w\in\mathfrak{F}^{\circ}$ the corresponding point in the zz–plane. For almost every $u\in[-L,L]$ we have
$\int_{0}^{L}\rho(u,v){\mathrm{d}}v\geq\mathrm{Im}\,\left(\mathcal{P}(\mathbf{p}^{-1}(u+iL))\right)~~\text{when }~~u\in U^{0}$ (4.8)
$\int_{0}^{L}\rho(u,v){\mathrm{d}}v\geq\mathrm{Im}\,\left(\mathcal{P}(\mathbf{p}^{-1}(u+iL))+\mathcal{P}(w)\right)~~\text{when }~~u\in U^{+},$ (4.9)
$\int_{0}^{L}\rho(u,v){\mathrm{d}}v\geq\mathrm{Im}\,\left(\mathcal{P}(\mathbf{p}^{-1}(u+iL))-\mathcal{P}(w)\right)~~\text{when }~~u\in U^{-}.$ (4.10)
Proof. Consider first the case where $U^{0}\ni u=\mathbf{p}(w)$, where $w\in{\mathbb{R}}$. The integral then computes the total variation of $\mathcal{P}$ along the orthogonal trajectory $\gamma_{w}^{w_{1}}:=\mathbf{p}^{-1}(u+i[0,L])$, which starts from the point $w=\mathbf{p}^{-1}(u)\in{\mathbb{R}}$ and ends at $w_{1}=\mathbf{p}^{-1}(u+iL)$:
$\int_{0}^{L}\rho(u,v){\mathrm{d}}v=\int_{0}^{L}\left|\frac{{\mathrm{d}}\mathcal{P}}{{\mathrm{d}}z}\right|\left|\frac{{\mathrm{d}}z}{{\mathrm{d}}\mathbf{p}}\right||{\mathrm{d}}\zeta|=\int_{\gamma_{w}^{w_{1}}}\left|\frac{{\mathrm{d}}\mathcal{P}}{{\mathrm{d}}z}\right||{\mathrm{d}}z|.$ (4.11)
Since $w\in{\mathbb{R}}$, we have in particular $\mathrm{Im}\,\mathcal{P}(w)=0$ , and we conclude that the total variation is certainly at least equal to $\mathrm{Im}\,\mathcal{P}(w_{1})=\mathrm{Im}\,\mathcal{P}(\mathbf{p}^{-1}(u+iL))$. This establishes the inequality (4.8) in this case.
Let us now assume that $u\in U^{+}$ so that $\mathcal{L}_{d}^{\perp}(w)\cup\mathcal{K}\neq\emptyset$. Let $x_{0}$ be the first intersection point (with $\mathcal{K}$) along $\mathcal{L}_{d}^{\perp}(w)$ as it leaves $w\in\mathfrak{F}$ and $x_{1}$ be the last such point (it may happen that $x_{0}=x_{1}$). Then the total variation of $\mathcal{P}$ can be bounded below by the sum of $\mathrm{Im}\,\mathcal{P}(\mathbf{p}^{-1}(u+iL))$ on $x_{1},\mathbf{p}^{-1}(u+iL)]$ and $\mathrm{Im}\,\mathcal{P}(w)$ on $w,x_{0}]$, which proves the second inequality (4.8) (note $\mathrm{Im}\,\mathcal{P}=0$ on $\mathcal{K}$).
Now consider a point $u\in U^{-}$. Assume the corresponding recessive orthogonal trajectory $\mathcal{L}^{\perp}(w)$ does not intersects $\mathcal{K}$. Then the total variation of $\mathcal{P}$ can be bounded below by the sum of $\mathrm{Im}\,\mathcal{P}(\mathbf{p}^{-1}(u+iL))-\mathrm{Im}\,\mathcal{P}(w)$ on $w,\mathbf{p}^{-1}(u+iL)]$. If the recessive trajectory $\mathcal{L}^{\perp}(w)$ does intersect $\mathcal{K}$, then we recover the second inequality in (4.8), which implies the third inequality since $\mathrm{Im}\,\mathcal{P}\geq 0$ in $\mathbb{H}$}}. $\blacksquare$
Lemma 4.4
Let $L$ be sufficiently large. Then for almost every $u\in[-L,L]$:
$\mathrm{Im}\,\mathcal{P}(\mathbf{p}^{-1}(u+iL))=L+\frac{({\mathcal{I}}(\mathfrak{F})-{\mathcal{I}}(\mathcal{K}))L}{(L^{2}+u^{2})}+\mathcal{O}(L^{-2}).$ (4.12)
Proof. Since
$\mathbf{p}(z)=z+\frac{{\mathcal{I}}(\mathfrak{F})}{z}+\mathcal{O}(z^{-2}),~\mathcal{P}(z)=z+\frac{{\mathcal{I}}(\mathcal{K})}{z}+\mathcal{O}(z^{-2}),$ (4.13)
we have
$\mathcal{P}(\mathbf{p}^{-1}(\zeta))=\zeta-\frac{{\mathcal{I}}(\mathfrak{F})-{\mathcal{I}}(\mathcal{K})}{\zeta}+\mathcal{O}(\zeta^{-2}).$ (4.14)
The statement is a consequence of (4.14). $\blacksquare$
Lemma 4.5
In the large $L$ limit the area $\mathscr{A}_{L}$ of $\mathscr{R}_{L}$ satisfies
$\mathscr{A}_{L}=\iint_{R_{L}}{\rho(u,v)}^{2}{\mathrm{d}}^{2}\zeta\geq 2L^{2}+\pi({\mathcal{I}}(\mathfrak{F})-{\mathcal{I}}(\mathcal{K}))+\mathcal{O}(L^{-1}).$ (4.15)
Proof. By the definition of $U^{\pm}$, there is a bijection $U^{+}\mapsto U^{-}$ given by $u\mapsto u^{\star}$. The pull back of $du$ from $U^{-}$ to $U^{+}$ yields
$\phi(u)du=\dfrac{\frac{\partial\mathrm{Im}\,\mathbf{p}}{\partial n_{-}}}{\frac{\partial\mathrm{Im}\,\mathbf{p}}{\partial n_{+}}}du\leq du,$ (4.16)
where $n_{+}$ is the normal vector to $\mathfrak{F}^{\circ}$ in the dominant direction. Now, according to Lemmas 4.3,4.4, we have
$\int_{-L}^{L}\int_{0}^{L}\rho(u,v){\mathrm{d}}v{\mathrm{d}}u\geq\int_{U^{+}}\mathrm{Im}\,\left(\mathcal{P}(\mathbf{p}^{-1}(u+iL))+\mathcal{P}(w)\right)du+\int_{U^{-}}\mathrm{Im}\,\left(\mathcal{P}(\mathbf{p}^{-1}(u+iL))-\mathcal{P}(w)\right)du$ (4.17)
$+\int_{U^{0}}\mathrm{Im}\,\mathcal{P}(\mathbf{p}^{-1}(u+iL)du=\int_{-L}^{L}\left(L+\frac{[{\mathcal{I}}(\mathfrak{F})-{\mathcal{I}}(\mathcal{K})]L}{(L^{2}+u^{2})}+\mathcal{O}(L^{-2})\right){\mathrm{d}}u+\int_{U^{+}}\mathrm{Im}\,\mathcal{P}(\mathbf{p}^{-1}(u)))\left(1-\phi(u)\right)du\geq$ (4.18)
$2L^{2}+\frac{\pi}{2}[{\mathcal{I}}(\mathfrak{F})-{\mathcal{I}}(\mathcal{K})]+\mathcal{O}(L^{-1}),$ (4.19)
where we used (4.16) and the fact that $U^{+}\cup U^{-}\cup U^{0}=[-L,L]$ up to a measure zero set.
We rewrite the latter inequality as follows
$\iint_{R_{L}}\Big(\rho(u,v)-1\Big){\mathrm{d}}^{2}\zeta\geq\frac{\pi}{2}[{\mathcal{I}}(\mathfrak{F})-{\mathcal{I}}(\mathcal{K})]+\mathcal{O}(L^{-1}).$ (4.20)
To finally estimate the area (which is the integral of the square of $\rho$) we proceed as follows:
$\mathscr{A}_{L}-2L^{2}=\iint_{R_{L}}\Big(\rho(u,v)^{2}-1\Big){\mathrm{d}}^{2}\zeta=\iint_{R_{L}}\Big(\rho(u,v)-1\Big)^{2}{\mathrm{d}}^{2}\zeta+2\iint_{R_{L}}\Big(\rho(u,v)-1\Big){\mathrm{d}}^{2}\zeta\geq$
[…]
The proof is complete. $\blacksquare$
Upper bound for (4.5).
Lemma 4.6
In the large $L$ limit the area $\mathscr{A}_{L}$ of $\mathscr{R}_{L}$ satisfies
$\mathscr{A}_{L}=2L^{2}+\mathcal{O}(L^{-1}).$ (4.22)
Proof. According to Prop. 3.6 the map $\xi=\mathcal{P}(z(\mathbf{p}))$ can be defined as a univalent mapping provided we perform the additional slits from the stagnation points as described in that proposition and that were denoted by $\Sigma$. Then from (4.3), noting that $\Sigma$ has zero measure,
[…]
Thus we can rewrite the integral in the $\xi$–plane ($\xi=\mathcal{P}(z)$) as a regular area integral with respect to the standard Lebesgue $\xi$–measure ${\mathrm{d}}^{2}\xi$:
$\mathscr{A}_{L}=\iint_{\mathscr{R}_{L}}{\mathrm{d}}^{2}\xi.$ (4.24)
We remind that the domain of integration $\mathscr{R}_{L}$ is the deformed rectangle, bounded by the $\mathcal{P}$-image of the outer boundary of ${\mathcal{R}}_{L}$, namely bounded by
$\gamma_{left}$ $=\mathcal{P}\left(\mathbf{p}^{-1}(-L+i[0,L])\right),\ \ \gamma_{top}=\mathcal{P}\left(\mathbf{p}^{-1}([-L,L]+iL)\right),$ (4.25)
$\gamma_{right}$ $=\mathcal{P}\left(\mathbf{p}^{-1}(L+i[0,L])\right),\ \ \gamma_{bottom}=[\mathcal{P}(\mathbf{p}(-L)),\mathcal{P}(\mathbf{p}(L)]\subset{\mathbb{R}}.$ (4.26)
The $\xi$–area of this region can be computed by means of Green formula
$\mathscr{A}_{L}=\frac{1}{2i}\oint_{\partial\mathscr{R}_{L}}\overline{\xi}{\mathrm{d}}\xi$ (4.27)
The computation of this line integral can be parametrized by $\zeta\in\partial R_{L}$ using the fact that $\xi=\mathcal{P}\circ\mathbf{p}^{-1}(\zeta)=\zeta-\frac{\Delta}{\zeta}+\mathcal{O}(\zeta^{-2})$. The integration on the segment of the real $\xi$–axis yields a zero contribution, and so we are left with the integral over the left, top and right sides of $\mathscr{R}_{L}$; in the $\zeta$ plane these are just the straight segments $\widetilde{\gamma}_{\text{left}}=-L+i[0,L]$, $\widetilde{\gamma}_{\text{top}}=iL+[-L,L]$ and $\widetilde{\gamma}_{\text{right}}=L+i[L,0]$, respectively. Expanding the terms we obtain
$\mathscr{A}_{L}=\frac{1}{2i}\int_{\widetilde{\gamma}_{\text{left,top,right}}}\left(\overline{\zeta}-\frac{\Delta}{\overline{\zeta}}+\mathcal{O}(\zeta^{-2})\right)\left(1+\frac{\Delta}{\zeta^{2}}+\mathcal{O}(\zeta^{-3})\right){\mathrm{d}}\zeta=$
$=2L^{2}+\frac{\Delta}{2i}\int_{\widetilde{\gamma}_{\text{left,top,right}}}\left(\frac{\overline{\zeta}}{\zeta^{2}}-\frac{1}{\overline{\zeta}}\right){\mathrm{d}}\zeta+\mathcal{O}(L^{-1}).$ (4.28)
At this point one has to explicitly compute the parametrized integral indicated above and verify that the result vanishes, leaving only the subleading corrections. We leave the calculus exercise to the reader. See also a similar computation on page 61 [16], below (4.14) ibidem. $\blacksquare$
Proof of Theorem 4.2.
Suppose that the compact (finite union of continua) $\mathcal{K}$ has Jenkins’ interception property with respect to $\mathfrak{F}$. Then from Lemmas 4.3-4.6 it follows that
$2L^{2}+\pi[{\mathcal{I}}(\mathfrak{F})-{\mathcal{I}}(\mathcal{K})]+\mathcal{O}(L^{-1})\leq\mathscr{A}_{L}=2L^{2}+\mathcal{O}(L^{-1}).$ (4.29)
Therefore
${\mathcal{I}}(\mathfrak{F})\leq{\mathcal{I}}(\mathcal{K}).$ (4.30)
It only remains to explain how the equality can be achieved. If ${\mathcal{I}}(\mathfrak{F})-{\mathcal{I}}(\mathcal{K})]=0$ then $\mathscr{A}_{L}$ is $2L^{2}+o(1)$. For this to happen in the chain of inequalities (4) we must also have
$\iint_{R_{L}}\Big(\rho(u,v)-1\Big)^{2}{\mathrm{d}}^{2}\zeta=0,$ (4.31)
namely, $\rho\equiv 1$ (up to sets of zero measure). This implies that for $\zeta$ sufficiently large, where $\mathcal{P}$ and $\mathbf{p}$ are both univalent, we have
$\left|\frac{{\mathrm{d}}\mathcal{P}}{{\mathrm{d}}\mathbf{p}}\right|\equiv 1$ (4.32)
so that the equality holds everywhere by analytic continuation. At this point this means that $\mathcal{P}(z)=c\mathbf{p}(z)+r$ with $|c|=1$ and $r\in{\mathbb{C}}$. Since both $\mathcal{P},\mathbf{p}$ map the real axis onto itself and the upper half-plane in the upper half–plane, we must have $c=1$ and $r\in{\mathbb{R}}$. From their asymptotic expansion for large $z$ it follows that $r=0$. Thus $\mathcal{P}\equiv\mathbf{p}$ and hence $\mathcal{K}=\mathfrak{F}$. $\blacksquare$
Immediate consequences of Theorem 4.2 are Corollary 4.7 and Theorem 4.10 stated below.
Corollary 4.7
Suppose that in the conditions of Theorem 4.2 for all $z$ in the relative interior of each arc of $\mathfrak{F}$ we have $\mathscr{L}^{\perp}(z)\cap\mathcal{K}\neq\emptyset$ for both orthogonal trajectories emanating from $z$. Then ${\mathcal{I}}(\mathfrak{F})\leq{\mathcal{I}}(\mathcal{K})$; moreover, the equality implies $\mathcal{K}=\mathfrak{F}$.
To state further consequences of Theorem 4.2, we remind/introduce some classes of poly-continua. Let $E=\{e_{1},...e_{N}\}\subset{{\mathbb{H}}}$ be a finite set of (distinct) anchor points. We denote by ${\mathbb{K}}_{E}$ the set of compacts $\mathcal{K}$ in $\mathbb{H}$}} such that
  • $\mathcal{K}$ is a finite union of continua, i.e., a poly-continuum;
  • $E\subset\mathcal{K}$, and every continuum of $\mathcal{K}$ connects two different points of $E$ or a point of $E$ with $\mathbb{R}$}.
Consider the case where all arcs in $\mathfrak{F}$ from Theorem 4.2 possess the $S$-property, namely, $\mathfrak{F}$ is a Zakharov-Shabat spectrum introduced in Section 2 and associated to a Boutroux quadratic differential $Q$. Then either of the two orthogonal trajectories in $\mathcal{L}^{\perp}(z)$, $z\in\mathfrak{F}^{\circ}$ can be considered dominant and we can modify Definition 4.1 as follows:
Definition 4.8
Suppose that $\mathfrak{F}$ has the SS–property (2.18) at all points of $\mathfrak{F}^{\circ}$. Then we say that a compact set $\mathcal{K}$ possesses Jenkins’ interception property (relative to $\mathfrak{F}$) if for every $z\in\mathfrak{F}$ we have $\mathcal{L}^{\perp}(z)\cap\mathcal{K}\neq\emptyset$ for at least one orthogonal trajectories emanating from $z$, possibly at $z$ itself. We denote the class of these sets by $\mathbb{K}_{\mathfrak{F}}$ .
Remark 4.9
It is this property that was used in [17].
Theorem 4.10
Let $\mathbb{K}_{\mathfrak{F}}$ denote the family in Definition 4.8. Then
${\mathcal{I}}(\mathfrak{F})=\min_{\mathcal{K}\in\mathbb{K}_{\mathfrak{F}}}{\mathcal{I}}(\mathcal{K})$ (4.33)
with the equality occurring if and only if $\mathcal{K}$ coincides with $\mathfrak{F}$.
Remark 4.11
Let us consider a set $E$ with $N=2$ anchor points $e_{1,2}$. Denote by $\mathfrak{F}$ the Zakharov–Shabat spectrum of the quasimomentum differential ${\mathrm{d}}\mathbf{p}_{2}$ on the RS $\mathfrak{R}_{2}$ branched at $E\cup\bar{E}$. Note that $\mathfrak{F}\in\mathbb{K}_{E}$ and for any $\mathcal{K}\in\mathbb{K}_{E}$ the points $e_{1},e_{2}$ belong to the same connected component of $\mathcal{K}\cup{\mathbb{R}}$. This implies $\mathbb{K}_{\mathfrak{F}}=\mathbb{K}_{E}$. Thus, by Theorem 4.10, $\mathfrak{F}$ is the minimizer of $\mathcal{I}(\mathcal{K})$ in $\mathbb{K}_{E}$.
We can now turn to the proof of Theorem 1.3.
4.1 Proof of Theorem 1.3
We consider $Q{\mathrm{d}}z^{2}$, a quasimomentum type quadratic differential with simple poles at the points $e_{j}\in E$, $j=1,\dots,N$. Let $\mathfrak{F}\in{\mathbb{K}}_{E}$ be the Zakharov–Shabat spectrum of $Q{\mathrm{d}}z^{2}$ and let $M=M(\mathfrak{F})$ be the connectivity matrix of $\mathfrak{F}$. We want to prove that $\mathfrak{F}$ is a minimizer of $\mathcal{I}(\mathcal{K})$ among the subclass ${\mathbb{K}}_{E,M}\subset{\mathbb{K}}_{E}$ that have the same or greater connectivity as $\mathfrak{F}$. According to Theorem 4.10, it is sufficient to prove that any $\mathcal{K}\in{\mathbb{K}}_{E,M}$ has Jenkins’ interception property (Definition 4.8) with respect to $\mathfrak{F}$.
Consider $z_{0}\in\mathfrak{F}^{\circ}$ and let $T$ be the connected component (continuum) of $\mathfrak{F}$ containing $z_{0}$. Without loss of generality we can assume that the two trajectories $\mathcal{L}^{\perp}(z_{0})$ emanating from $z_{0}$ split the plane $\mathbb{C}$} into two connected components with only one of them, called $G_{2}$, being adjacent to $\mathbb{R}$}. Indeed, there are only finitely many zeros of $Q$ and, thus, only finitely many $z\in\mathfrak{F}^{\circ}$, such that any of the trajectories $\mathcal{L}^{\perp}(z)$ contains one of these zeros. By construction, there is a point $e_{j}\in T\cap G_{1}\cap E$. If $T$ is connected with $\mathbb{R}$}, then $\mathcal{K}$ also contains a continuum $\mathcal{K}_{0}$ connecting $e_{j}$ and $\mathbb{R}$}. Thus, $\mathcal{K}_{0}$ must intersect the boundary of $G_{1}$, i.e., at least one of the trajectories $\mathcal{L}^{\perp}(z_{0})$ intersects $\mathcal{K}$. If $T$ is not connected with $\mathbb{R}$}, then there is another point $e_{k}\in T\cap G_{2}\cap E$ and a continuum $\mathcal{K}_{l}\subset\mathcal{K}$ containing both $e_{j},e_{k}$. Then $K_{l}$ must intersect $\partial G_{1}$ and we again prove that at least one of the trajectories $\mathcal{L}^{\perp}(z_{0})$ intersects $\mathcal{K}$. Thus, $\mathcal{K}$ has the Jenkins’ Interception property relative to $\mathfrak{F}$. Now the statement follows from Theorem 4.10. $\blacksquare$
We complete this section with last corollary of Theorem 4.10. Consider the RS $\mathfrak{R}_{N}$ defined by the branchpoints $E\cup\bar{E}$. Then the Zakharov–Shabat spectrum $\mathfrak{F}\in{\mathbb{K}}_{E}$ of the quasimomentum differential ${\mathrm{d}}\mathbf{p}_{N}$ on $\mathfrak{R}_{N}$ minimizes $\mathcal{I}(\mathcal{K})$ among all possible (Schwarz symmetrical) branchcuts of $\mathfrak{R}_{N}$. Before formulating this result we want to describe such collections of branchcuts.
Define the subclass $\mathbb{L}_{E}\subset{\mathbb{K}}_{E}$ that consists of arcs connecting anchor points $E$ with each other and with $\mathbb{R}$} in such a way that each $e_{j}$ has an odd number of emanating arcs and there are no closed loops, i.e., ${{\mathbb{H}}}\setminus\mathcal{K}$ is connected. In other words, class $\mathbb{L}_{E}$ consists of all possible Schwarz symmetrical branchcuts for $\mathfrak{R}_{N}$. It is clear that $\mathfrak{F}\in\mathbb{L}_{E}$.
Corollary 4.12
The unique minimizer $\mathfrak{F}$ of the Dirichlet energy in the class $\mathbb{L}_{E}\subset{\mathbb{K}}_{E}$ is the Zakharov–Shabat spectrum of the real normalized quasimomentum differential ${\mathrm{d}}\mathbf{p}_{N}$ on $\mathfrak{R}_{N}$, i.e., Zakharov–Shabat spectrum is the minimizing poly-continuum among all possible branchcuts of $\mathfrak{R}_{N}$:
${\mathcal{I}}(\mathfrak{F})=\min_{\mathcal{K}\in\mathbb{L}_{E}}{\mathcal{I}}(\mathcal{K}),$ (4.34)
with the equality occurring if and only if $\mathcal{K}$ coincides with $\mathfrak{F}$.
Proof. We start with a brief discussion about trees and valences of their vertices. Any system of branch-cuts $\mathfrak{C}$ for the Riemann surface $\cal R_{N}$ is such that $E\cup\overline{E}$ are odd-valent vertices and all other vertices, if any, are necessarily even-valent.
We are only considering systems of cuts for which the complement is connected, so that $\mathfrak{C}$ must be a forest (union of trees). Since the sum of the degrees of all vertices in a tree is twice the number of edges, there must be an even number of odd-valent vertices in any tree.
In particular this implies that, for any interior point $p$ on any edge $e$ of a tree $T$, the number of odd–valent vertices in each component of $T\setminus\{p\}$ is odd. Indeed, if we split any edge $e$ at a point $p$ within a tree $T$ we obtain two trees $T_{1},T_{2}$, each with one added vertex $p$ of valence $1$ on the edge $e$. Since each $T_{1},T_{2}$ must have an even number of odd-valent vertices, including the added one, we conclude that the number of the original odd-valent vertices in each $T_{1},T_{2}$ is odd.
Our arguments now are similar to the proof of Theorem 1.3 presented above. Let $\mathcal{K}\in\mathbb{L}_{E}$ be a set of branch-cuts for $\cal R_{N}$ branched at $E\cup\overline{E}$. Let $z_{0}\in\mathfrak{F}$ be in relative interior of $\mathfrak{F}$ and let the regions $G_{1},G_{2}$ be defined as in the proof of Theorem 1.3 above. We have seen that $E\cap G_{1}$ has an odd number of points, from which we deduce that not all connected components of $\mathcal{K}$ that intersect $G_{1}$ can be entirely contained in $G_{1}$. Indeed, each such component can only contain an even number of points of $E$. Thus $\mathcal{K}$ must intersect the boundary of $G_{1}$ and hence at least one of $\mathcal{L}^{\perp}(z_{0})$. Now the statement follows from Theorem 4.10. $\blacksquare$
5 Hausdorff continuity of the Dirichlet energy
For the purpose of this section we consider a more general polynomial external field for the weighted Greens’ energy and, correspondingly, the following, more general, Dirichlet problem (see Problem 3.1).
Let us fix $r\in{\mathbb{N}}$ and $r$ real parameters $t_{1},t_{2},\dots,t_{r}$. Denote by $\Phi$ the polynomial external field
$\Phi(z):=\sum_{\ell=1}^{r}t_{\ell}z^{\ell},\ \ t_{r}>0.$ (5.1)
Problem 5.1
For a given poly-continuum $\mathcal{K}\subset\overline{{{\mathbb{H}}}}$, let $G$ be the solution of the Dirichlet problem
  1. 1.
    $G$ is continuous and bounded on $\mathbb{H}$}};
  2. 2.
    $G$ is harmonic outside $\mathcal{K}$;
  3. 3.
    $G$ satisfies the boundary conditions $G(z)=\mathrm{Im}\,\Phi(z),\ \ \forall z\in\mathcal{K}\cup{\mathbb{R}}$.
Recall (see after Problem 3.1) that poly-continua are regular for the above problem as well. The condition $t_{r}>0$ is only for definiteness; if $t_{r}<0$ we can re-map the problem to an equivalent one for which $t_{r}>0$ by swapping the upper with the lower half plane and $\Phi(z)\mapsto-\overline{\Phi(\overline{z})}$.
Since the external field is harmonic, there is a real signed measure ${\mathrm{d}}\rho_{\mathcal{K}}$ supported on the outer boundary of $\mathcal{K}$ such that
$G(z)=\int_{\partial\mathcal{K}}\ln\left|\frac{z-\overline{w}}{z-w}\right|{\mathrm{d}}\rho_{\mathcal{K}}(w)$ (5.2)
Similarly to Section 3, we will denote by $\mathcal{P}$ the following function analytic on the universal cover of ${{\mathbb{H}}}\setminus\mathcal{K}$;
$\mathcal{P}(z)=\mathcal{P}(z;\mathcal{K})=\Phi(z)-iG(z)+H(z)=\Phi(z)-g(z)$ (5.3)
where $H(z)$ is the harmonic conjugate function of $G$ and g(z):=iG(z)H(z)g(z):=iG(z)-H(z). The following properties are simply ascertained (see (3.9)).
Proposition 5.2
The function $\mathcal{P}(z)=U(z)+iV(z)$ satisfies that $V(z)$ is zero on $\mathcal{K}\cup{\mathbb{R}}$, continuous in $\mathbb{H}$}} and harmonic in ${{\mathbb{H}}}\setminus\mathcal{K}$. Moreover we have, for $|z|\to\infty$
$\mathcal{P}(z)=\sum_{\ell=1}^{r}t_{\ell}z^{\ell}+\sum_{\ell\geq 1}\frac{\mathcal{I}_{\ell}(\mathcal{K})}{\ell z^{\ell}},$ (5.4)
where
$\mathcal{I}_{\ell}=\mathcal{I}_{\ell}(\mathcal{K})=2\int_{\partial\mathcal{K}}{\mathrm{Im}\,(w^{\ell}){\mathrm{d}}\rho_{\mathcal{K}}(w)}.$ (5.5)
and here ${\mathrm{d}}\rho_{\mathcal{K}}$ is the positive measure in the Poisson-Jensen representation (3.1) and supported on the outer boundary of $\mathcal{K}$.
The Dirichlet energy $D_{\mathcal{K},\Phi}$ of $G$ is defined by any of the following formulae
$\pi D_{\mathcal{K},\Phi}=\pi\iint_{H^{+}}|{\rm grad}\,G|^{2}{\mathrm{d}}^{2}z=2\ \iint_{\partial\mathcal{K}\times\partial\mathcal{K}}\ln\left|\frac{z-\overline{w}}{z-w}\right|{\mathrm{d}}\rho_{\mathcal{K}}(w){\mathrm{d}}\rho_{\mathcal{K}}(z)=2\int_{\mathcal{K}}\mathrm{Im}\,\Phi(z){\mathrm{d}}\rho_{\mathcal{K}}(z).$ (5.6)
These equalities imply that each of the expressions is strictly positive, which can be written also as the positivity of the expression
$\mathcal{I}_{\Phi}(\mathcal{K}):=2\int_{\mathcal{K}}\mathrm{Im}\,\Phi(z){\mathrm{d}}\rho_{\mathcal{K}}(z)=\sum_{\ell=1}^{r}t_{\ell}\mathcal{I}_{\ell}(\mathcal{K})=-\mathop{\mathrm{res}}_{z=\infty}\Phi(z){\mathrm{d}}\mathcal{P}(z)>0.$ (5.7)
where the last residue is computed by extending $\Phi,\mathcal{P}$ to the whole plane $\mathbb{C}$} by Schwartz-symmetry. In the case $\Phi(z)=z$, we keep our previous notation $\mathcal{I}_{\Phi}(\mathcal{K})=\mathcal{I}(\mathcal{K})$.
Let $\mathcal{K}_{1,2}$ denote compact sets in $\mathbb{H}$$\cup{\mathbb{R}}$ and let ${\mathrm{d}}_{H}$ denote the Hausdorff metric between compact sets:
${\mathrm{d}}_{H}(\mathcal{K}_{1},\mathcal{K}_{2})=\max\left\{\sup_{x\in\mathcal{K}_{1}}{\rm dist}(x,\mathcal{K}_{2}),\sup_{y\in\mathcal{K}_{2}}{\rm dist}(y,\mathcal{K}_{1})\right\}.$ (5.8)
If ${\mathrm{d}}_{H}(\mathcal{K}_{1},\mathcal{K}_{2})=\epsilon$ then we have that $K_{1}\subset K_{2}^{\epsilon}$ and viceversa, where $\mathcal{K}^{\epsilon}$ is the $\epsilon$–fattening of a set
$\mathcal{K}^{\epsilon}=\bigcup_{x\in\mathcal{K}}\mathbb{D}_{\epsilon}(x).$ (5.9)
Let $Q_{{}_{\mathcal{K}}}(x;y)$ denote the Green function of the complement of the poly-continuum $\mathcal{K}$ in $\mathbb{H}$}}, namely:
  1. 1.
    $\forall y\not\in\mathcal{K}\cup{\mathbb{R}}$ the function $h_{{}_{\mathcal{K}}}(x;y):=Q_{{}_{\mathcal{K}}}(x;y)-\ln\left|\frac{x-\overline{y}}{x-y}\right|$ extends to a harmonic function of $x$ in a neihbourhood of $y$; it is harmonic and bounded in ${{\mathbb{H}}}\setminus\mathcal{K}$ and continuous in $\mathbb{H}$}};
  2. 2.
    $Q_{\mathcal{K}}(x;y)$ vanishes identically for $x\in\mathcal{K}\cup{\mathbb{R}}$.
Since $\mathcal{K}$ is a poly-continuum, it has no component of zero capacity. Viceversa we could rephrase the Dirichlet problem quasi-everywhere (i.e. up to sets of zero capacity). There is no practical advantage in one formulation versus the other and we stick to the above one.
We start with a useful definition.
Definition 5.3
A compact $\mathcal{K}$ will be called “Dirichlet regular” if all its connected components have logarithmic capacity not less than some $s>0$. A family $\mathscr{K}$ of compact sets is “uniformly Dirichlet regular” if the infimum of all capacities of all the connected components of each $\mathcal{K}\in\mathscr{K}$ is greater than zero.
Proposition 5.4
Let $\mathcal{K}$ be a Dirichlet regular (Definition 5.3) compact set so that the Green function for the domain ${{\mathbb{H}}}\setminus\mathcal{K}$ is well defined and continuous up to the boundary. Let $\mathcal{C}$ be any compact set with positive distance from $\mathcal{K}$. If $\mathcal{K}_{n}$ is a sequence of uniformly Dirichlet regular compact sets that converges to $\mathcal{K}$ in Hausdorff topology, then
$\sup_{(x,y)\in{{\mathbb{H}}}\times\mathcal{C}}\left|Q_{{}_{\mathcal{K}_{n}}}(x;y)-Q_{{}_{\mathcal{K}}}(x;y)\right|\to 0.$ (5.10)
In particular the Green functions converge uniformly to each other in any closed set at finite distance from $\mathcal{K}$.
Proof. Let $\mathcal{K}_{1},\mathcal{K}_{2}$ be two Dirichlet regular compact sets with ${\mathrm{d}}_{H}(\mathcal{K}_{1},\mathcal{K}_{2})\leq\epsilon$. Let $\mathcal{C}$ be another closed set without intersection with either one. Then Corollary A.5 implies that there is a constant $d_{0}>0$ such that
$Q_{{}_{\mathcal{K}_{j}}}(x;y)\leq d_{0}\sqrt{{\rm dist}(x,\mathcal{K}_{j})},\ \ \ \ \forall y\in\mathcal{C}.$ (5.11)
Now, using the maximum principle for harmonic functions, we have that for all $x\in{{\mathbb{H}}}$ and $y\in\mathcal{C}$
$Q_{{}_{\mathcal{K}_{1}}}(x;y)-\max_{\bullet\in\mathcal{K}_{2}}Q_{{}_{\mathcal{K}_{1}}}(\bullet;y)\leq Q_{{}_{\mathcal{K}_{2}}}(x;y)$ (5.12)
$Q_{{}_{\mathcal{K}_{1}}}(x;y)-\max_{\bullet\in\mathcal{K}_{2}\atop w\in\mathcal{C}}Q_{{}_{\mathcal{K}_{1}}}(\bullet;w)\leq Q_{{}_{\mathcal{K}_{2}}}(x;y)\ \ \ \Rightarrow$
$Q_{{}_{\mathcal{K}_{1}}}(x;y)-Q_{{}_{\mathcal{K}_{2}}}(x;y)\leq\max_{\bullet\in\mathcal{K}_{2}\atop w\in\mathcal{C}}Q_{{}_{\mathcal{K}_{1}}}(\bullet;w).$ (5.13)
Swapping the roles of $\mathcal{K}_{1}\leftrightarrow\mathcal{K}_{2}$ we have also
$Q_{{}_{\mathcal{K}_{2}}}(x;y)-Q_{{}_{\mathcal{K}_{1}}}(x;y)\leq\max_{\bullet\in\mathcal{K}_{1}\atop w\in\mathcal{C}}Q_{{}_{\mathcal{K}_{2}}}(\bullet;w).$ (5.14)
Since $\mathcal{K}_{1}$ is in the $\epsilon$ neighbourhood of $\mathcal{K}_{2}$, and viceversa, from (5.11) follows that
$\max\left\{\max_{\bullet\in\mathcal{K}_{2}\atop w\in\mathcal{C}}Q_{{}_{\mathcal{K}_{1}}}(\bullet;w),\max_{\bullet\in\mathcal{K}_{1}\atop w\in\mathcal{C}}Q_{{}_{\mathcal{K}_{2}}}(\bullet;w)\right\}\leq d_{0}\sqrt{\epsilon}$ (5.15)
and thus we have
$\big|Q_{{}_{\mathcal{K}_{1}}}(x;y)-Q_{{}_{\mathcal{K}_{2}}}(x;y)\big|\leq d_{0}\sqrt{\epsilon}\ \ \ \ \ \ \forall(x,y)\in{{\mathbb{H}}}\times\mathcal{C}.$ (5.16)
Let now $\mathcal{K}_{n}$ be a sequence converging to $\mathcal{K}$ with ${\mathrm{d}}_{H}(\mathcal{K},\mathcal{K}_{n})<\epsilon_{n}\to 0$. We have assumed that it is uniformly Dirichlet regular. By the Corollary A.5 we then have that there is $d_{0}>0$ (depending on $\mathcal{C}$) such that the estimate (5.16) holds uniformly for the sequence:
$\big|Q_{{}_{\mathcal{K}}}(x;y)-Q_{{}_{\mathcal{K}_{n}}}(x;y)\big|\leq d_{0}\sqrt{\epsilon_{n}}\ \ \ \ \ \ \forall(x,y)\in{{\mathbb{H}}}\times\mathcal{C}.$ (5.17)
The proof follows immediately. $\blacksquare$
We want to use Proposition 5.4 to prove the continuity of the Dirichlet energies for our external field $\phi$. To this end we have the following Proposition.
Proposition 5.5
The energy $\mathcal{I}_{\Phi}(\mathcal{K})$ for the Problem 5.1 is given by
$\mathcal{I}_{\Phi}(\mathcal{K})=\oint_{|z|,|w|>>1}\Phi(z)\Phi(w)\partial_{z}\partial_{w}Q_{{}_{\mathcal{K}}}(z;w)\frac{{\mathrm{d}}z{\mathrm{d}}w}{2\pi^{2}}.$ (5.18)
In this formula the Green function is extended to $\mathbb{C}$$\times{\mathbb{C}}$ by Schwartz-symmetry.
Proof. If $G_{{}_{\mathcal{K}\cup\overline{\mathcal{K}}}}(z;w)$ is the Green function for ${\mathbb{C}}\setminus\mathcal{K}\cup\overline{\mathcal{K}}$ then
$Q_{{}_{\mathcal{K}}}(z;w)=G_{{}_{\mathcal{K}\cup\overline{\mathcal{K}}}}(z;w)-G_{{}_{\mathcal{K}\cup\overline{\mathcal{K}}}}(z;\overline{w}),$ (5.19)
and this formula extends to the whole complement of $\mathcal{K}\cup\overline{\mathcal{K}}$ in $\mathbb{C}$} in such a way that
$Q_{{}_{\mathcal{K}}}(z;w)=-Q_{{}_{\mathcal{K}}}(z;\overline{w})=-Q_{{}_{\mathcal{K}}}(\overline{z};w).$ (5.20)
Note that the differential $\partial_{w}Q(z,w){\mathrm{d}}w$ is meromorphic for $w\not\in\mathcal{K}$, see (2.11), with a simple pole at $w=z,\overline{z}$ of residues $\frac{1}{2},\frac{1}{2}$, respectively. It is also harmonic w.r.t. $z$ and zero for $z\in\mathcal{K}$. Take a (union of) closed contour(s) separating $z,\overline{z}$ from $\mathcal{K}\cup\overline{\mathcal{K}}$, with $z,\overline{z}$ in the exterior and $\mathcal{K}\cup\overline{\mathcal{K}}$ in the interior, and consider the expression
$G(z)=\mathrm{Im}\,\left[\oint_{w\in\gamma}\Phi(w)\partial_{w}Q(z,w)\frac{{\mathrm{d}}w}{2i\pi}\right].$ (5.21)
By the residue theorem, this is the same as (using that $\mathrm{Im}\,{\Phi(z)}=-\mathrm{Im}\,(\Phi(\overline{z}))$)
$G(z)=\mathrm{Im}\,\left[\Phi(z)+\oint_{|w|=R>|z|}\Phi(w)\partial_{w}Q(z,w)\frac{{\mathrm{d}}w}{2i\pi}\right].$ (5.22)
The formula (5.21) shows that $G(z)$ is bounded for $z\not\in\mathcal{K}$, while the formula (5.22) shows that $G$ tends to $\mathrm{Im}\,\Phi$ on the boundary of $\mathcal{K}$ thanks to the fact that $\mathcal{K}$ is a poly-continuum and hence regular for the Dirichlet problem (see comment after Problem 5.1). Thus we have established that $G(z)$ is indeed the solution of the Dirichlet Problem 5.1. In particular we have
$\mathrm{Im}\,(\mathcal{P})=\mathrm{Im}\,\left[i\oint_{|w|>>1}\!\!\!\!\!\Phi(w)\partial_{w}Q(z,w)\frac{{\mathrm{d}}w}{2\pi}\right]\ \ \ \Rightarrow\ \ \ {\mathrm{d}}\mathcal{P}=2\partial_{z}\mathrm{Im}\,(\mathcal{P}){\mathrm{d}}z=i\oint_{|w|>>1}\!\!\!\!\!\Phi(w)\partial_{z}\partial_{w}Q(z,w)\frac{{\mathrm{d}}w}{\pi}$
We now use the residue expression in formula (5.7)
$\mathcal{I}_{\Phi}(\mathcal{K})=\oint_{|z|>>1}\Phi(z)\frac{{\mathrm{d}}\mathcal{P}(z)}{2i\pi}=\oint_{|z|>>1,|w|=2|z|}\!\!\!\!\!\Phi(z)\Phi(w)\partial_{z}\partial_{w}Q_{{}_{\mathcal{K}}}(z,w)\frac{{\mathrm{d}}w{\mathrm{d}}z}{2\pi^{2}}.$ (5.23)
This proves the statement. $\blacksquare$
Theorem 5.6
Let $E=\{e_{1},\dots,e_{N}\}\subset{{\mathbb{H}}}$ be a finite set of (pairwise distinct) points. For a fixed $R>0$, let $\mathbb{K}_{R}\subset{\mathbb{K}}_{E}$ be the subclass of poly-continua $\mathcal{K}$ contained in the disk $|z|\leq R$. For a fixed polynomial $\phi=\mathrm{Im}\,\left(\sum_{\ell=1}^{r}{t_{\ell}}z^{\ell}\right)$, the map $\mathcal{I}_{\Phi}(\mathcal{K}):\mathbb{K}_{R}\mapsto{\mathbb{R}}$, where $\mathcal{I}_{\Phi}(\mathcal{K})$ is given by (5.7), is continuous in Hausdorff topology.
Proof. The set $\mathbb{K}_{R}$ is closed in Hausdorff topology, a simple exercise. It also follows that this class is uniformly Dirichlet regular: indeed if $x,\widetilde{x}\in E$ are any two distinct points that belong to the same connected component of $\mathcal{K}\in\mathbb{K}_{R}$, then the capacity of that component is at least 4|xx~|4|x-$\widetilde${x}|. Thus the class $\mathbb{K}_{R}$ is also uniformly Dirichlet regular in the sense of Definition 5.3.
Then the proof follows from Propositions 5.4, 5.5, which prove uniform convergence of the harmonic functions $Q_{\mathcal{K}}$ (and hence also of their derivatives) over the contours of integration in (5.23). $\blacksquare$
6 Existence of the minimizer in ${\mathbb{K}}_{E,M}$
We now revert to the original problem with the external field given by $\Phi=z$. The goal of this section is to prove that, for any set of anchors $E$ and connectivity matrix $M$, the Dirichlet energy $\mathcal{I}(\mathcal{K})$ attains its minimum in the class ${\mathbb{K}}_{E,M}$. In view of Theorem 5.6 all that we need is to show that there exists a fixed rectangle $R_{E}\subset{{\mathbb{H}}}\cup{\mathbb{R}}$ and a minimizing sequence $\{\mathcal{K}_{n}\}\subset{\mathbb{K}}_{E,M}$, such that $\mathcal{K}_{n}\subset R_{E}$ for all $n\in{\mathbb{N}}$.
We start with the observation that if $\{\mathcal{K}_{n}\}\subset\mathbb{K}_{E,M}$ is a minimizing sequence, i.e., the sequence $\mathcal{I}_{n}=\mathcal{I}(\mathcal{K}_{n})$ converges to $\breve{I}=\inf_{\mathcal{K}\in{\mathbb{K}}_{E,M}}\{\mathcal{I}(\mathcal{K})\}$, then there exists a minimizing sequence $\{\hat{\mathcal{K}}_{n}\}\subset\mathbb{K}_{E,M}$, where each $\hat{\mathcal{K}}_{n}$ is a collection of piecewise smooth contours. Indeed, let us consider closed $\varepsilon_{n}$ fattening $\mathring{\mathcal{K}}_{n}\subset\mathbb{K}_{E,M}$ of each $\mathcal{K}_{n}$, where $\varepsilon_{n}>0$ is so small that $|\mathcal{I}(\mathring{\mathcal{K}}_{n})-\mathcal{I}(\mathcal{K}_{n})|<\frac{1}{n}$. The later inequality follows from the continuity of the energy functional $\mathcal{I}(\mathcal{K})$ on $\mathbb{K}_{E,M}\cap R_{E}$, see Theorem 5.6. Thus, $\{\mathring{\mathcal{K}}_{n}\}$ is also a minimizing sequence. Now, in each closed domain $\mathring{\mathcal{K}}_{n}$ we choose a piecewise smooth contour $\hat{\mathcal{K}}_{n}\subset\mathring{\mathcal{K}}_{n}$ connecting points of $E\subset\mathring{\mathcal{K}}_{n}$ between themselves and with $\mathbb{R}$} according to the given connectivity $M$. Thus, $\hat{\mathcal{K}}_{n}\subset\mathbb{K}_{E,M}$. But, according to the Jenkins interception property, see Theorem 4.2, $\mathcal{I}(\mathring{\mathcal{K}}_{n})>\mathcal{I}(\hat{\mathcal{K}}_{n}).$ Thus, $\hat{\mathcal{K}}_{n}\subset\mathbb{K}_{E,M}$ is also a minimizing sequence, where each $\hat{\mathcal{K}}_{n}$ consists of piecewise smooth contours.
Our approach consists of two parts: first proving that for any minimizing sequence $\{\mathcal{K}_{n}\}\subset{\mathbb{K}}_{E,M}$ there exists $b>0$ such that $K_{n}$ lies in the horizontal strip $0\leq\mathrm{Im}\,z\leq b$ for all $n\in{\mathbb{N}}$ and, secondly, proving that there exists $a>0$, such that the rectangle $R_{E}$, bounded by $|\mathrm{Re}\,z|\leq a$, $0\leq\mathrm{Im}\,z\leq b$, contains a minimizing sequence. Without loss of generality we can assume that minimizing sequences considered below consist of piecewise smooth poly-continua.
Lemma 6.1
For any minimizing sequence $\{\mathcal{K}_{n}\}\subset\ {\mathbb{K}}_{E,M}$ there exists $b>0$ such that $\mathcal{K}_{n}$ lies in the horizontal strip $0\leq\mathrm{Im}\,z\leq b$ for all $n\in{\mathbb{N}}$.
Proof. Let $z_{n}\in\mathcal{K}_{n}$ is such that $\mathrm{Im}\,z_{n}=\max_{z\in\mathcal{K}_{n}}\{\mathrm{Im}\,z\}$. Assume, to the contrary, that no horizontal strip contains all the poly-continua $\mathcal{K}_{n}$, $n\in{\mathbb{N}}$. Then $\mathrm{Im}\,z_{n}\to\infty$ as $n\to\infty$. Set $\tilde{\mathcal{K}}_{n}=\mathcal{K}_{n}\cap\left\{|z-z_{n}|\leq\frac{1}{2}\right\}$. Then $\tilde{\mathcal{K}}_{n}$ contains a component connecting $z_{n}\in\tilde{\mathcal{K}}_{n}$ with $\{|z-z_{n}|=\frac{1}{2}\}\cap\mathcal{K}$. We observe that the logarithmic capacity ${\rm cap}(\tilde{\mathcal{K}}_{n})$ of $\tilde{\mathcal{K}}_{n}$ is at least $\frac{1}{8}$.
Let $\mu_{n}$ denotes the total mass one positive equilibrium Borel measure on $\tilde{\mathcal{K}}_{n}$ (with respect to free logarithmic energy.) Recalling the Green energy functional $J_{0}$ (1.3), note that the Dirichlet energy $\mathcal{I}(\mathcal{K}_{n})=-2J_{0}[\rho(\mathcal{K}_{n})]\geq-2J_{0}[\mu_{n}]$ , where $\rho(\mathcal{K}_{n})$ is the (positive) measure minimizing the Green energy, i.e. $J_{0}[\rho(\mathcal{K}_{n})]=\mathfrak{J}(\mathcal{K}_{n})$, see (3.7). Note that
$-4\int_{\tilde{\mathcal{K}}_{n}}\mathrm{Im}\,zd\mu_{n}(z)\leq-2(2\mathrm{Im}\,z_{n}-1)\to$-$\infty$ (6.1)
as $\mathrm{Im}\,z_{n}\to\infty$, where the left hand side is the linear part of the Green energy (1.3) with $\mathcal{K}=\tilde{\mathcal{K}}_{n}$ and $\rho_{\mathcal{K}}=\mu_{n}$. We now represent the remaining quadratic term of $J_{0}(\mu_{n})$ as
$2\iint_{\tilde{\mathcal{K}}_{n}\times\tilde{\mathcal{K}}_{n}}\ln\left|\frac{z-\overline{w}}{z-w}\right|{\mathrm{d}}\mu_{n}(z){\mathrm{d}}\mu_{n}(w)=$~~~~~~~~~~~~~~~~~~~ (6.2)
$2\iint_{\tilde{\mathcal{K}}_{n}\times\tilde{\mathcal{K}}_{n}}\ln\left|z-\overline{w}\right|{\mathrm{d}}\mu_{n}(z){\mathrm{d}}\mu_{n}(w)-2\iint_{\tilde{\mathcal{K}}_{n}\times\tilde{\mathcal{K}}_{n}}\ln\left|z-w\right|{\mathrm{d}}\mu_{n}(z){\mathrm{d}}\mu_{n}(w).$ (6.3)
The first term of the last sum behaves like $O(\ln\mathrm{Im}\,z_{n})$ as $n\to\infty$, The second term is bounded by $2|\ln{\rm cap}(\tilde{\mathcal{K}}_{n})|$. Thus, in view of (6.1), we conclude that $J_{0}(\rho(\mathcal{K}_{n}))\leq J_{0}(\mu_{n})$ approaches -$\infty$ as $n\to\infty$. Thus, $\{\mathcal{K}_{n}\}$ cannot be a minimizing sequence for the Dirichlet energy as $\mathcal{I}(\mathcal{K}_{n})\to\infty$ as $n\to\infty$. $\blacksquare$
According to Lemma 6.1, a minimizing sequence $\{\mathcal{K}_{n}\}$ must be contained within a strip $S=\{z\in{\mathbb{C}}:\,0\leq\mathrm{Im}\,z\leq b\}$ for some $b>0$. The next lemma show that if the poly-continua $\mathcal{K}_{n}\in\ {\mathbb{K}}_{E,M}$, $n\in{\mathbb{N}}$, protrude in $S$ towards infinity, say, on the left, then for any $a\leq\min_{j}\mathrm{Re}\,e_{j}$ we can construct another poly-continua $\tilde{\mathcal{K}}_{n}\in\ {\mathbb{K}}_{E,M}$ satisfying $a=\min_{z\in\tilde{\mathcal{K}}_{n}}\mathrm{Re}\,z$ and such that $\mathcal{I}_{\mathcal{K}_{n}}>\mathcal{I}_{\tilde{\mathcal{K}}_{n}}$, thus constructing a minimizing sequence $\{\tilde{\mathcal{K}}_{n}\}$ in a semistrip (bounded from the left). Repeating one more time the same arguments we can construct a minimizing sequence contained in the rectangle $R_{E}$.
To formulate the result, we go back to the notion of the generalized quasimomentum (3.8).
If a smooth oriented arc $\gamma$ is part of $\mathcal{K}$, surrounded by its complement ${{\mathbb{H}}}\setminus\mathcal{K}$, then it is well known ([32]) that
$\frac{\partial}{\partial n_{+}}V_{+}+\frac{\partial}{\partial n_{-}}V_{-}=2\pi u$ (6.4)
on $\gamma$, where $u(z)$, $z\in\gamma$, is the density of the equilibrium measure on $\mathcal{K}$ and $n_{\pm}$ denote the positive/negative unit normal on $\gamma$.
Denote by $\tilde{u}(z)$ the average of the boundary values of $\mathcal{P}$ on $\gamma$, i.e.,
$-\tilde{u}(z)=\frac{1}{2}\left[\mathcal{P}_{+}+\mathcal{P}_{-}\right]=\frac{1}{2}\left[U_{+}+U_{-}\right],$ (6.5)
since $V_{\pm}=0$ on $\gamma$. By the Cauchy-Riemann equations, we get
$\frac{\ \partial}{\ \partial\zeta}U_{\pm}=\pm\frac{\ \partial}{\ \partial n_{\pm}}V.$ (6.6)
where $\zeta$ denotes the arclength parameter on $\gamma$. It follows then from (6.5) that
$\frac{\ \partial}{\ \partial n_{+}}V_{+}-\frac{\ \partial}{\ \partial n_{-}}V_{-}=-2\frac{\ \partial}{\ \partial\zeta}\tilde{u}$, (6.7)
that is, $2\frac{\ \partial}{\ \partial\zeta}\tilde{u}$ represents the mismatch of normal derivatives of $V$ (we remind that the S-property corresponds to the zero mismatch). Combining this equation with (6.4), we get
$\frac{\ \partial}{\ \partial n_{+}}V_{+}=\pi u-\frac{\ \partial}{\ \partial\zeta}\tilde{u},\qquad\frac{\ \partial}{\ \partial n_{-}}V_{-}=\pi u+\frac{\ \partial}{\ \partial\zeta}\tilde{u}$, (6.8)
According to (6.7) the side of dominant orthogonal trajectory $\mathcal{L}^{\perp}_{d}(z)$ when $z\in\gamma$, is determined by the sign of $\frac{d}{d\zeta}\tilde{u}(\zeta)$, see (6.5). Suppose $\gamma$ is a vertical segment oriented upwards. Then ${\mathrm{d}}z=i{\mathrm{d}}\zeta$, and from (3.8) we get
$\frac{d}{d\zeta}\left\{\left[U_{+}(z)+iV_{+}(z)\right]+\left[U_{-}(z)+iV_{-}(z)\right]\right\}=i(\mathcal{P}_{+}^{\prime}(z)+\mathcal{P}_{-}^{\prime}(z))=\int\mathrm{Re}\,\left[\frac{1}{z-\bar{w}}-\frac{1}{z-w}\right]u(w)|dw|$ (6.9)
i.e.,
$-2\frac{d}{d\zeta}\tilde{u}(z)=\int\mathrm{Re}\,\left[\frac{1}{z-\bar{w}}-\frac{1}{z-w}\right]u(w)|dw|.$ (6.10)
Let $\mathcal{K}\in{\mathbb{K}}_{E,M}$ and $\hat{\mathcal{K}}=\mathcal{K}\cap\{z:\mathrm{Re}\,z\leq a\}\neq\emptyset$ for some $a$ such that $a\leq\min_{j}\mathrm{Re}\,e_{j}$. Let $l=\{z:\mathrm{Re}\,z=a,~\mathrm{Im}\,z\geq 0\}$. We want to construct $\tilde{\mathcal{K}}\in{\mathbb{K}_{{E,M}}}$, such that $\tilde{\mathcal{K}}$ coincides with $\mathcal{K}$ for all $\mathrm{Re}\,z>a$ and $\tilde{K}\cap\{z:\mathrm{Re}\,z<a\}=\emptyset$. Below are the steps of how we construct $\tilde{\mathcal{K}}$ from $\mathcal{K}$, that is, how we define $\tilde{\mathcal{K}}\cap l$. For any connected component $S$ of $\hat{K}$ such that $S\cap{\mathbb{R}}\neq\emptyset$, replace $S\cap l$ by a segment $[a,a+ib]$, where $b=\max\{\mathrm{Im}\,z:~\mathrm{Re}\,z=a~{\rm and}~z\in S\}$. If the supremum of all $b$ constructed above is $B=\max\{\mathrm{Im}\,z:~\mathrm{Re}\,z=a~{\rm and}~z\in\hat{\mathcal{K}}\}$, we denote
$\tilde{\mathcal{K}}=\mathcal{K}\cap\{z:\mathrm{Re}\,z\geq a\}\cup[a,a+iB].$ (6.11)
Otherwise, let $\sup b=B_{1}<B$. Now, for any connected component $S$ of $^{\hat{\mathcal{K}}}$ such that $S\cap{\mathbb{R}}=\emptyset$ and $S$ intersects $[a,a+iB]$ at at least two points, replace $S\cap l$ by $b_{1},b_{2}]$, where $b_{1}$ is the infinum and $b_{2}$ is the supremum of $\mathrm{Im}\,z$ among those points of intersection. If a connected component $S$ of $^{\hat{\mathcal{K}}}$ intersects $[a,a+iB]$ only at one point (and $S\cap{\mathbb{R}}=\emptyset$), we add this point to $\tilde{\mathcal{K}}$.
Lemma 6.2
[…]
Proof. The proof is based on applying the Jenkins’ interception property of Theorem 4.2. From the construction of $\tilde{\mathcal{K}}$ it is sufficient to prove that the dominant orthogonal trajectories on the boundary $\mathrm{Re}\,z=a$ of $\tilde{\mathcal{K}}$ are on the positive side, i.e., go to the left since the boundary is oriented upwards. That is, let a segment $\gamma$ on $\mathrm{Re}\,z=a$ belong to $\tilde{\mathcal{K}}$ and let $z\in\gamma$ be an interior point. We need to show that $\mathcal{L}_{d}^{\perp}(z)$ is directed to the left. If that is true, then $\mathcal{L}_{d}^{\perp}(z)\cap\mathcal{K}\neq\emptyset$ by the construction of $\tilde{\mathcal{K}}$. So, in view of (6.7), (6.5) and (6.10), we need to show that the integral in (6.10) is non negative. This follows from the inequality below, because $\mathrm{Re}\,z\leq\mathrm{Re}\,w$ for any $w\in\tilde{\mathcal{K}}$,
$\mathrm{Re}\,\left[\frac{1}{z-\bar{w}}-\frac{1}{z-w}\right]=\mathrm{Re}\,(z-w)\left[\frac{1}{|z-\bar{w}|^{2}}-\frac{1}{|z-w|^{2}}\right]\geq 0$ (6.12)
and since $|z-w|^{2}\leq|z-\bar{w}|^{2}$. $\blacksquare$
Corollary 6.3
Given any anchor set $E\subset{{\mathbb{H}}}$ and arbitrary connectivity matrix $M$, there exists a uniformly bounded sequence $\{\mathcal{K}_{n}\}\subset{\mathbb{K}}_{E,M}$ minimizing the Dirichlet energy $\mathcal{I}(\mathcal{K})$ in ${\mathbb{K}}_{E,M}$.
Proof. Since $\mathcal{I}(\mathcal{K})\geq 0$, a minimizing sequence $\{\mathcal{K}_{n}\}$ for $\mathcal{I}$ exists. According to Lemma 6.1, all the poly-continua from $\{\mathcal{K}_{n}\}$ must be located in a horizontal strip of $\mathbb{H}$}}, adjacent to $\mathbb{R}$}. Then Lemma 6.2, asserts that if minimizing poly-continua $\{\mathcal{K}_{n}\}$ are unbounded, there exists another minimizing sequence $\{\tilde{\mathcal{K}}_{n}\}\subset{\mathbb{K}}_{E,M}$ that consists of uniformly bounded poly-continua. $\blacksquare$
Since the Dirichlet energy $\mathcal{I}(\mathcal{K})$ is Hausdorff continuous on a set $K_{E,M}$, where all the poly-continua are uniformly bounded, and since such set is closed in the Hausdorff topology, we obtain the following theorem, which implies Theorem 1.1.
Theorem 6.4
For any anchor set $E\subset{{\mathbb{H}}}$ and for any connectivity matrix $M$ there exists a poly-continuum $\mathcal{K}$ minimizing the Dirichlet energy $\mathcal{I}(\mathcal{K})$ within ${\mathbb{K}}_{E,M}$.
The Theorem 1.2 is addressed in Section 7.
7 Schiffer variations and S-curves
Assume that the positive Borel measure ${\mathrm{d}}\mu_{\mathcal{K}}$ minimizes the Greens’ energy functional $J_{0}$ in (1.3) (see (3.7)), where ${\rm supp\,}\mu\subset\mathcal{K}$. For given set of anchors $E=\{e_{1},...e_{N}\}\subset{{\mathbb{H}}}$ define $E(z):=\prod_{j=1}^{N}(z-e_{j})(z-\overline{e}_{j})$. Suppose that the poly-continuum $\mathcal{K}\in{\mathbb{K}}_{E}$ is critical so that the variation of the energy (1.3) is zero. For given smooth and bounded $h:\bar{{\mathbb{H}}}\to{\mathbb{C}}$, the Schiffer variation represents the infinitesimal variation of the energy under the action of the (infinitesimal) diffeomorphism generated by the vector field $\dot{z}=h(z)$. We are interested only in the diffeomorphisms of the upper half plane that fix the set of anchors, and hence $h(e)=0$ for $e\in E$ and $h(x)\in{\mathbb{R}}$ for $x\in{\mathbb{R}}$. The variation formula is given by
$0=\mathrm{Re}\,\left(\iint\left(\frac{h(z)-\overline{h}(w)}{z-\overline{w}}-\frac{h(z)-h(w)}{z-w}-\right){\mathrm{d}}\mu(z){\mathrm{d}}\mu(w)+2\int\Phi^{\prime}(z)h(z){\mathrm{d}}\mu(z)\right)$ (7.1)
with $\Phi(z)=iz$ (or more generally any germ of analytic function in the upper half plane such that $\mathrm{Re}\,\Phi$ is defined on the support of ${\mathrm{d}}\mu$, single–valued, and zero on $\mathbb{R}$}) see, for example, [26], Section 3. For the formula (7.1) to be correct it is actually sufficient that $h(z)$ is bounded in a neighbourhood of the support of ${\mathrm{d}}\mu$.
Let $x\in{\mathbb{C}}$: we derive the following formulas for $x$ outside of the support of ${\mathrm{d}}\mu$ and then explain how they are actually still valid within the support. We use the Schiffer condition (7.1) with two different choices of $h(z)$ (which we denote $h(z),k(z)$) as follows:
$h(z)$ $=E(z)\left[\frac{1}{z-x}+\frac{1}{z-\overline{x}}\right]=\frac{E(x)}{z-x}+\frac{E(\overline{x})}{z-\overline{x}}+\frac{E(z)-E(x)}{z-x}+\frac{E(z)-E(\overline{x})}{z-\overline{x}}=$ (7.2)
$=\frac{E(x)}{z-x}+\frac{E(\overline{x})}{z-\overline{x}}+Q(z,x)+Q(z,\overline{x}).$ (7.3)
Observe that $Q(z,x)$ is a polynomial with real coefficients in $z,x$ of degree 2N12N-1 defined by the middle expression above. Similarly we define
$k(z)=iE(z)\left[\frac{1}{z-x}-\frac{1}{z-\overline{x}}\right]=\frac{iE(x)}{z-x}-\frac{iE(\overline{x})}{z-\overline{x}}+iQ(z,x)-iQ(z,\overline{x}).$ (7.4)
Observe that both $h,k$ defined above vanish for $z\in E$ and are real (or zero) for $z\in{\mathbb{R}}$, as requested. Plugging $h(z)$ into (7.1) and simplifying we obtain
[…]
$-(\overline{w}\leftrightarrow w)\bigg){\mathrm{d}}\mu(z){\mathrm{d}}\mu(w)+2\int\Phi^{\prime}(z)E(z)\left[\frac{1}{z-x}-\frac{1}{z-\overline{x}}\right]{\mathrm{d}}\mu(z)\Bigg]=0$ (7.5)
Define
$S(z,\overline{w},x):=\frac{Q(z,x)-Q(\overline{w},x)}{z-\overline{w}}.$ (7.6)
Then we can rewrite (7.5) as follows
$\mathrm{Re}\,\Bigg[$ $\iint\left(\frac{-E(x)}{(z-x)(\overline{w}-x)}+\frac{-E(\overline{x})}{(z-\overline{x})(\overline{w}-\overline{x})}+S(z,\overline{w},x)+S(z,\overline{w},\overline{x})-(\overline{w}\leftrightarrow w)\right){\mathrm{d}}\mu(z){\mathrm{d}}\mu(w)+$
$+2\Phi^{\prime}(x)E(x)\int\frac{{\mathrm{d}}\mu(z)}{z-x}+2\int\frac{\big(\Phi^{\prime}(z)E(z)-\Phi^{\prime}(x)E(x)\big){\mathrm{d}}\mu(z)}{z-x}-(x\leftrightarrow\overline{x})\Bigg]=0.$ (7.7)
We assume that $\overline{\Phi(\overline{z})}=-\Phi(z)$ so that we can conjugate all the terms containing $\overline{x}$ in (7.7) and obtain (recalling that $S(z,\overline{w},x)$ is a polynomial with real coefficients):
$\mathrm{Re}\,\Bigg[$ $\iint\left(\frac{-E(x)}{(z-x)(\overline{w}-x)}+\frac{-E(x)}{(\overline{z}-x)(w-x)}+S(z,\overline{w},x)+S(\overline{z},w,x)-(\overline{w}\leftrightarrow w)\right){\mathrm{d}}\mu(z){\mathrm{d}}\mu(w)+$
$+2\Phi^{\prime}(x)E(x)\int\left[\frac{1}{z-x}-\frac{1}{\overline{z}-x}\right]{\mathrm{d}}\mu(z)+$
$+2\int\left(\frac{\big(\Phi^{\prime}(z)E(z)-\Phi^{\prime}(x)E(x)\big)}{z-x}+\frac{\big({\Phi^{\prime}(\overline{z})}E(\overline{z})-{\Phi^{\prime}(x)}E(x)\big)}{\overline{z}-x}\right){\mathrm{d}}\mu(z)\Bigg]=$
$=\mathrm{Re}\,\Bigg[$ $E(x)\iint\left(\frac{1}{\overline{z}-x}-\frac{1}{z-x}\right)\left(\frac{1}{\overline{w}-x}-\frac{1}{w-x}\right){\mathrm{d}}\mu(z){\mathrm{d}}\mu(w)+$
$-2\Phi^{\prime}(x)E(x)\int\left[\frac{1}{\overline{z}-x}-\frac{1}{z-x}\right]{\mathrm{d}}\mu(z)+\mathscr{S}(x)+\mathscr{Q}(x)\Bigg]$ (7.8)
where
$\mathscr{S}(x):=$ $\iint\bigg[S(z,\overline{w},x)+S(\overline{z},w,x)-S(z,w,x)-S(z,\overline{w},x)\bigg]{\mathrm{d}}\mu(z){\mathrm{d}}\mu(w)$ (7.9)
$\mathscr{Q}(x):=$ $2\int\left(\frac{\big(\Phi^{\prime}(z)E(z)-\Phi^{\prime}(x)E(x)\big)}{z-x}+\frac{\big(\Phi^{\prime}(\overline{z})E(\overline{z})-\Phi^{\prime}(x)E(x)\big)}{\overline{z}-x}\right){\mathrm{d}}\mu(z)$ (7.10)
We observe that $\mathscr{S}$ is a polynomial with real coefficients in $x$ of degree $\leq 2N-1$. Viceversa $\mathscr{Q}(x)$ is a real–analytic function of $x$ with the same singularities as $\Phi^{\prime}(x)$. In the case $\Phi$ is a polynomial of degree $r$, see (5.1), $\mathscr{Q}(x)$ is a polynomial of degree at most 2N+r22N+r-2. If we define
$g^{\prime}(x):=i\int\left[\frac{1}{z-x}-\frac{1}{\overline{z}-x}\right]{\mathrm{d}}\mu(z)=i\partial_{x}\int\ln\left(\frac{x-\overline{z}}{x-z}\right){\mathrm{d}}\mu(z),$ (7.11)
the equation (7) becomes:
$\mathrm{Re}\,\left[-E(x)\left(g^{\prime}(x)\right)^{2}-2i\Phi^{\prime}(x)E(x)g^{\prime}(x)+\mathscr{S}(x)+\mathscr{Q}(x)\right]=0.$ (7.12)
We now address the issue of what happens when $x$ is in the support. To prove the equation for (almost) any $x\in{\mathbb{C}}$, one has to repeat the steps of Lemma 5.1 from [26]. In the case when the equilibrium measure $\mu=\mu_{\mathcal{K}}$ is positive, for example, when $\Phi(z)=z$, these steps are literary the same. In the case when $\Phi$ is a more general analytic function (e.g. a polynomial) the measure $\mu=\mu_{\mathcal{K}}$ can become a signed measure. In this case the monotonically increasing function $m(r)$ from the proof of Lemma 5.1 becomes a function of bounded variation (BV), which is the difference of two monotonic functions. It is known that a BV function admits derivative almost everywhere and therefore, the approach of Lemma 5.1 is still valid in the case of signed measures, i.e., in the case of the general $\Phi(z)$ considered here.
Now we repeat the whole computation with $k(z)$ in (7.4). Using the same steps one verifies that the final equation is
$\mathrm{Im}\,\left[-E(x)\left(g^{\prime}(x)\right)^{2}-2i\Phi^{\prime}(x)E(x)g^{\prime}(x)+\mathscr{S}(x)+\mathscr{Q}(x)\right]=0.$ (7.13)
The equations (7.12), (7.13) together imply the following identity:
$E(x)\left(g^{\prime}(x)\right)^{2}+2i\Phi^{\prime}(x)E(x)g^{\prime}(x)=\mathscr{S}(x)+\mathscr{Q}(x)\ \ \ \Leftrightarrow\$(7.14)
$E(x)\Big(g^{\prime}(x)+i\Phi^{\prime}(x)\Big)^{2}=-E(x)\Phi^{\prime}(x)^{2}+\mathscr{S}(x)+\mathscr{Q}(x).$ (7.15)
We have proved:
Proposition 7.1
Suppose $\Phi(z)$ is a polynomial of degree $r$ with real coefficients and $E(x)=\prod_{j=1}^{N}(x-e_{j})(x-\overline{e}_{j})$. The complexified Green potential $g(x)$ of a Schiffer critical measure $\mu=\mu_{\mathcal{K}}$ satisfies
$\Big(g^{\prime}(x)+i\Phi^{\prime}(x)\Big)^{2}=-\Phi^{\prime}(x)^{2}+\frac{\mathscr{S}(x)+\mathscr{Q}(x)}{E(x)},$ (7.16)
where $\mathscr{S}(x),\mathscr{Q}(x)$ are polynomials with real coefficients of degree not exceeding 2N12N-1 and 2N+r22N+r-2 respectively.
Equation (7.16) implies that $i{\mathrm{d}}g=\sqrt{Q(z)}{\mathrm{d}}z$, where $Q{\mathrm{d}}z^{2}$ is a Boutroux quadratic differential. In the case when $\Phi=iz$ (so that $r(=\deg\Phi)=1$), $Q{\mathrm{d}}z^{2}$ is the quasimomentum type quadratic differential from main Theorem 1.2. It then follows that ${\rm supp}\,\mu_{\mathcal{K}}$ is the Zakharov–Shabat spectrum of $Q{\mathrm{d}}z^{2}$. (This statement also follows from [26], Lemma 5.2.) This completes the proof of Theorem 1.2.
Appendix A Green functions
In this appendix we collect some useful, and probably well-known facts about Green functions that are used in the main text. Nonetheless we could not find a direct and explicit reference in the literature about these properties and we think some readers may find them of independent interest.
We first establish the desired property of the ordinary Green function in the plane, and then transfer those statement to analogous statements for Green functions in the upper half-plane by a simple application of the “reflection principle”.
Theorem A.1
Let $\mathcal{K}$ be a continuum, $C={\rm Cap}(\mathcal{K})$ its capacity. Let $\Omega$ denote the unbounded connected component of the complement $\mathcal{K}^{c}$ and $G_{\Omega}(z,w)$ its Green function. Then for all $z,w\in{\mathbb{C}}$ we have
$G_{\Omega}(z;w)\leq\sqrt{\frac{{\rm dist}(z,\mathcal{K})}{C\,{\rm dist}(w,\mathcal{K})|z-w|}}{\rm e}^{U(w)}$ (A.1)
where
$U(w)=G(w;\infty)+\ln C=\int_{\mathcal{K}}\ln|w-t|{\mathrm{d}}\mu(t)$ (A.2)
and ${\mathrm{d}}\mu(t)$ is the harmonic (probability) measure on $\partial\Omega$.
Remark A.2
The reason the inequality is written in terms of the logarithmic potential $U$ rather than the Green potential $G$ is to make it apparent the role played by the capacity $C$. For a domain of zero capacity the inequality is vacuous.
Proof. To simplify the arguments without loss of generality we assume that the complement of $\mathcal{K}$ is connected (and unbounded). Let $\Omega=\mathcal{K}^{c}$ and $\Phi:\Omega\to\mathbb{D}^{c}$ be the uniformizing map to the exterior of the unit disk. We denote by $z=F(\zeta)={\zeta}{C}+\mathcal{O}(1)$ the inverse, $F:\mathbb{D}^{c}\to\Omega$. Here $C={\rm Cap}(\mathcal{K})$ is the capacity of $\mathcal{K}$. For brevity we shall also denote $\zeta=\Phi(z),\ \eta=\Phi(w)$ below. The function $C^{-1}F(\zeta)$ satisfies the assumptions of Lemma A.3 and thus (A.14) reads
${\rm dist}(z,\mathcal{K})\geq C\frac{(|\zeta|-1)^{2}}{|\zeta|}=4C\,\sinh\left(\frac{G(z;\infty)}{2}\right)^{2}\geq C\,G(z;\infty)^{2},$ (A.3)
where we note that $G(z;\infty)=\ln|\Phi(z)|$ is the Green function of $\Omega$. Thus we have the first inequality
$G(z;\infty)\leq\sqrt{\frac{{\rm dist}(z,\mathcal{K})}{C}}$ (A.4)
Now consider
$\widetilde{\Phi}(z):=\widetilde{\zeta}:=\frac{1-\zeta\overline{\eta}}{\zeta-\eta},\ \ \ \zeta=\frac{1+\widetilde{\zeta}\eta}{\widetilde{\zeta}+\overline{\eta}}=\eta+\frac{1-|\eta|^{2}}{\widetilde{\zeta}}+\mathcal{O}(\widetilde{\zeta}^{-2}).$ (A.5)
This maps $\Omega$ to the outside of $\mathbb{D}$, and $w$ to $\infty$. Define the function
$\widetilde{z}=\widetilde{F}(\widetilde{\zeta})=\frac{1}{F(\zeta)-F(\eta)}=\frac{\widetilde{\zeta}}{(1-|\eta|^{2})F^{\prime}(\eta)}+\mathcal{O}(1).$ (A.6)
which maps $\widetilde{\zeta}\in\mathbb{D}^{c}$ to $\widetilde{\Omega}=T(\Omega)$ and $\widetilde{z}=T(z)=\frac{1}{z-w}$. The previous inequality applied to $|\widetilde{\zeta}|={\rm e}^{G(z;w)}$ and with $\widetilde{C}:=\frac{1}{(|\eta|^{2}-1)|F^{\prime}(\eta)|}$ implies that
$G(z;w)=\ln|\widetilde{\zeta}|\leq\sqrt{\frac{{\rm dist}(\widetilde{z},\widetilde{\mathcal{K}})}{\widetilde{C}}}$ (A.7)
In the book [12], formula (21) on page 117 we read the following inequality:
$1-\frac{1}{|\zeta|^{2}}\leq\frac{|F^{\prime}(\zeta)|}{C}\leq\frac{|\zeta|^{2}}{|\zeta|^{2}-1},$ (A.8)
where the denominator is due to the fact that the quoted inequality is established in that reference for a normalized univalent function that behaves like $\zeta+\mathcal{O}(1)$ as $|\zeta|\to\infty$. Now (A.8) implies
$\frac{1}{\widetilde{C}}=(|\eta|^{2}-1){|F^{\prime}(\eta)|}\leq C{|\eta|^{2}}\ \ \ \Rightarrow\ \ \ \frac{1}{\widetilde{C}}\leq C{{\rm e}^{2G(w;\infty)}}=\frac{1}{C}{\rm e}^{2U(w)},$ (A.9)
where we have used the definition $U(w)=G(w;\infty)-\ln C$. On the other hand
${\rm dist}(\widetilde{z},\widetilde{\mathcal{K}})=\min_{t\in\mathcal{K}}\left|\frac{1}{z-w}-\frac{1}{t-w}\right|=\frac{1}{|z-w|}\min_{t\in\mathcal{K}}\left|\frac{z-t}{t-w}\right|.$ (A.10)
Now for all $t\in\mathcal{K}$ we have
$\min_{\bullet\in\mathcal{K}}\left|\frac{z-\bullet}{\bullet-w}\right|\leq\left|\frac{z-t}{t-w}\right|\leq\frac{|z-t|}{{\rm dist}(w,\mathcal{K})}.$ (A.11)
Since this is valid for all $t\in\mathcal{K}$ we can pass to the $\inf$ and get
${\rm dist}(\widetilde{z},\widetilde{\mathcal{K}})\leq\frac{{\rm dist}(z,\mathcal{K})}{|z-w|{\rm dist}(w,\mathcal{K})}.$ (A.12)
Plugging (A.12) and (A.9) into (A.7) gives
$G(z;w)\leq\sqrt{\frac{C\,{\rm dist}(z,\mathcal{K})}{{\rm dist}(w,\mathcal{K})|z-w|}}{\rm e}^{G(w;\infty)}=\sqrt{\frac{{\rm dist}(z,\mathcal{K})}{C\,{\rm dist}(w,\mathcal{K})|z-w|}}{\rm e}^{U(w)},$ (A.13)
which completes the proof. $\blacksquare$
Lemma A.3
Let $F(\zeta)=\zeta+a_{0}+\mathcal{O}(\zeta^{-1})$ be analytic and univalent for $|\zeta|>1$ and continuous for $|\zeta|\geq 1$. Then
$\delta:=\min_{|\rho|=1}|F(\zeta)-F(\rho)|\geq\frac{(|\zeta|-1)^{2}}{|\zeta|}.$ (A.14)
Viceversa we have
$|\zeta|-1\leq\frac{\delta}{2}+\sqrt{\frac{\delta^{2}}{4}+\delta}.$ (A.15)
Proof. Let $z_{0}\in\mathcal{K}^{c}$ and $w_{\star}\in\partial\mathcal{K}$ the (a) closest point; this point is accessible (see [30], pag. 277) because the straight segment from $z_{0}$ to $w_{\star}$ is inside the domain $\Omega:=F(\{|\zeta|\geq 1\})$. Consider then such a straight segment from $z_{0}$ to $w_{\star}$
$z(t)=z_{0}\,\frac{t}{\delta}+\left(1-\frac{t}{\delta}\right)w_{\star},\ \ \ t\in[0,\delta],\ \ \delta={\rm dist}(z_{0},\mathcal{K}),\quad t\in[0,\delta].$ (A.16)
Note that $|\dot{z}(t)|=1$. We try first to establish what is the maximum variation of $|\zeta(t)|=|\Phi(z(t))|$ along $z(t)$; namely we try to bound from below the distance from $\zeta\to\mathbb{D}$ as an increasing function of $\delta$, which gives then a lower bound of $\delta$ as a function of $|\zeta|$. We have
[…]
For the function $r(t):=|\zeta(t)|$ this inequality is saturated by the function
$r_{0}(t)-1=\frac{t}{2}+\sqrt{\frac{t^{2}}{4}+t}.$ (A.18)
Thus $r(t)\leq r_{0}(t)$ and $r=|\zeta|\leq r_{0}(\delta)$, namely
$|\zeta|\leq\frac{\delta}{2}+1+\sqrt{\frac{\delta^{2}}{4}+\delta}.$ (A.19)
Inverting the relation we have the statement (A.14). $\blacksquare$
Theorem A.4
Let $\mathcal{K}$ be an arbitrary compact set, $\mathcal{K}=\bigsqcup\mathcal{K}_{\mu}$ . Let $c$ be the minimum amongst the capacities of the components $\mathcal{K}_{\mu}$. Let $G$ be the Green function of the unbounded component $\Omega$ of ${\mathbb{C}}\setminus\mathcal{K}$, and similarly $G_{\mu}$ those for $\Omega_{\mu}$. Then for all $z,w\in{\mathbb{C}}$ we have
$G(z;w)\leq\sqrt{\frac{{\rm dist}(z,\mathcal{K})}{c\,{\rm dist}(w,\mathcal{K})|z-w|}}\min_{\mu}{\rm e}^{U_{\mu}(w)}.$ (A.20)
In particular, if $\mathcal{C}$ is a compact with finite distance from $\mathcal{K}$, we have that there is a constant $D$, depending on $\mathcal{C}$ such that
$G(z;w)\leq D\sqrt{\frac{{\rm dist}(z,\mathcal{K})}{c|z-w|}},\ \ \ \forall(z,w)\in{\mathbb{C}}\times\mathcal{C}.$ (A.21)
Proof. The formula (A.21) follows from (A.20) by taking he minimum of the continuous factors in $w$. Now, clearly $G(z,w)\leq G_{\mu}(z;w)$ for all $\mu$, so that taking the minimum over $\mu$ yields easily the claim. $\blacksquare$
We need an analogous property for the Green function of subsets of $\mathbb{H}$}}. Let $\mathcal{K}\subset{{\mathbb{H}}}$ be compact. Let us denote by $Q_{{}_{\mathcal{K}}}(z;w)$ the Green function of ${{\mathbb{H}}}\setminus\mathcal{K}$, namely
  1. 1.
    for any $w\not\in\mathcal{K}$ the function $z\mapsto Q_{{}_{\mathcal{K}}}(z;w)-\ln\left|\frac{z-\overline{w}}{z-w}\right|$ is harmonic in ${{\mathbb{H}}}\setminus\mathcal{K}$ and continuous in $\mathbb{H}$}};
  2. 2.
    $Q_{{}_{\mathcal{K}}}(z;w)\equiv 0$ for $w\not\in\mathcal{K}$ and $z\in\mathcal{K}\cup{\mathbb{R}}$.
It is a simple verification that
$Q_{{}_{\mathcal{K}}}(z;w)=G_{{}_{\mathcal{K}\cup\overline{\mathcal{K}}}}(z;w)-G_{{}_{\mathcal{K}\cup\overline{\mathcal{K}}}}(z;\overline{w})$ (A.22)
where $G_{{}_{\mathcal{K}}}$ is the ordinary Green function of ${\mathbb{C}}\setminus\mathcal{K}$. We then immediately have a similar Lemma
Corollary A.5
Let $\mathcal{K}\subset{{\mathbb{H}}}$ be a Dirichlet regular compact set, $c$ the minimal capacity of its components and $\mathcal{C}$ anoter disjoint compact. Then there is a constant $S>0$ such that
$Q_{{}_{\mathcal{K}}}(z;w)\leq S\sqrt{\frac{{\rm dist}(z,\mathcal{K})}{c}},\ \ \forall w\in\mathcal{C},\ \ z\in{{\mathbb{H}}}$ (A.23)
The same applies to a uniformly Dirichlet regular family.
Appendix B Maximal connectivity case: relation with Kuzmina’s Jenkins-Strebel differentials
The direct analogue of the well known Chebotarev problem corresponds to Theorem 1.1 with the ”maximal connectivity matrix” $M_{i,j}=1$, namely, the class consisting of continua containing all anchors points as well as a point on $\mathbb{R}$}. We denote such a matrix $M=\mathbf{1}$. Thus we have following weighted analogue of the Chebotarev problem.
Problem B.1
Given $E=\{e_{1},\dots,e_{N}\}\subset{{\mathbb{H}}}$, consider the class, $\mathbb{K}_{E,\mathbf{1}}$, consisting of continua (connected compact set) $\mathcal{K}$ such that $E\subset\mathcal{K}\subset\overline{{{\mathbb{H}}}}$ and $\mathcal{K}\cup{\mathbb{R}}$ is a connected set (equivalently, ${\mathrm{Out}}(\mathcal{K})$ is simply connected). The problem is to find a set $\mathcal{K}_{0}$ minimizing ${\mathcal{I}}(\mathcal{K})$ within $\mathbb{K}_{E,\mathbf{1}}$, and showing its uniqueness.
Theorem 1.3 implies that any other continuum in $\mathbb{K}_{E,\mathbf{1}}$ has the Jenkins Interception property relative to a $\mathfrak{F}\in\mathbb{K}_{E,\mathbf{1}}$, provided that such $\mathfrak{F}$ exists. In particular this implies the uniqueness of the minimum in the connectivity class of $\mathfrak{F}$. The existence is guaranteed by Theorem 1.1 which is, in a sense, the hardest part. But in this case we want to point out that the existence can be derived from existing theorems.
Thus, to completely address Problem B.1 we prove the existence of a Zakharov–Shabat spectrum within the class $\mathbb{K}_{E}^{(0)}=\mathbb{K}_{E,\mathbf{1}}$.
Proposition B.2
For any anchor set $E=\{e_{1},\dots,e_{N}\}\subset{{\mathbb{H}}}$ there is a (necessarily unique by the part (c) of Theorem 1.1) Zakharov–Shabat spectrum $\mathfrak{F}$ in the class $\mathbb{K}_{E}^{(0)}$ of Problem B.1, namely, such that ${{\mathbb{H}}}\setminus\mathfrak{F}$ is simply connected.
Proof. Kuzmina proved in [22] (see corrections in [23], in particular Theorem 3 ibidem) a theorem about the existence and uniqueness of a meromorphic quadratic differential with prescribed number of annular domains and disk domains in correspondence with second order poles with negative biresidues. If we specialize the theorem and require no annular domains and no double poles then Kuzmina’s theorem guarantees that the critical graph is connected and the complement in $\mathbb{P}^{1}$ is a union of domains of type half-plane; the half planes abut poles of higher orders with prescribed singular expansion.
Let us review Kuzmina’s theorem in a simplified formulation that is more amenable to our application. Let ${\bf A}=\{a_{1},\dots,a_{K}\}$. Fix $T_{0}\in{\mathbb{C}}$ and $T_{1}\in i{\mathbb{R}}$. Then there exists a unique quadratic differential $Q(z)$ with a pole of order $4$ at $z=\infty$ and with at most simple poles at $\bf A$}, such that
$Q(z)=\left((T_{0})^{2}+\frac{T_{1}T_{0}}{z}+\mathcal{O}(z^{-2})\right){\mathrm{d}}z^{2}$ (B.1)
and with connected critical graph. Then the complement of the graph in $\mathbb{C}$} is a union of two domains of type half plane (hence simply connected) abutting the point $\infty$ along the critical directions $\arg(T_{0}z)\in\{0,\pi\}$. Taking the square root of (B.1) we have
$\sqrt{Q}=\left(T_{0}+\frac{T_{1}}{2z}+\mathcal{O}(z^{-2})\right){\mathrm{d}}z$ (B.2)
so that we identify $T_{1}$ with the residue of the quadratic differential. For our application we need to set $T_{0}=1$, $T_{1}=0$, and ${\bf A}=E\cup\overline{E}$; in this case the uniqueness stated in the theorem implies that the critical graph $\Gamma$ is invariant under conjugation, $\Gamma=\overline{\Gamma}$. In particular the real axis must belong to the critical graph because there must be always one critical arc extending to the pole along the critical diractions (which in our case are $\arg(z)=0,\pi$). This proves the existence of the Zakharov–Shabat spectrum with simply connected complement in $\mathbb{H}$}}. $\blacksquare$
Remark B.3 (Bound on the genus of the surface)
The maximum genus of the double cover associated to the Zakharov–Shabat quadratic differential in this case coincides with the case when, for appropriate configurations, we have a unique double zero of $Q$ on $\mathbb{R}$} and all other simple zeros in $\mathbb{H}$}}; a count in this case shows that there are a maximum of 2N+2N22N+2N-2 branch-points so that the genus is 2N22N-2. See Fig. 1.
Conflict of Interest declaration.
The authors have no financial or proprietary interests in any material discussed in this article.
Data availability.
The software producing the pictures was coded by M. B. and is available at [2].